### v1

# RoHOI: Robustness Benchmark for Human-Object Interaction Detection

Abstract
Human-Object Interaction (HOI) detection is crucial for robot-human assistance, enabling context-aware support. However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate prediction. To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges. Despite advances, current models struggle with environmental variability, occlusion, and noise. Our benchmark, RoHOI, includes 20 corruption types based on HICO-DET and V-COCO datasets and a new robustness-focused metric. We systematically analyze existing models in the related field, revealing significant performance drops under corruptions. To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, dynamically adjusting the model’s optimization to enhance robust feature learning. Extensive experiments show our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection. Benchmarks, datasets, and code will be made publicly available at https://github.com/Kratos-Wen/RoHOI.
1 Introduction
Human-Object Interaction (HOI) detection is a key task in visual scene understanding, involving the identification and interpretation of interactions between humans and objects in an image [25, 19, 78, 56, 26, 20]. It serves as a foundation for many critical applications, including autonomous driving [45, 48], robotics [65], video surveillance [10, 50], and augmented reality [76], where a fine-grained understanding of human behavior and object interactions is paramount [13].
Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups, i.e., Optical system-induced artifacts, Sensor, compression and transmission artifacts, Environmentally induced artifacts, and Geometric and scene distortions. Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions. These corruptions simulate real-world degradation factors encountered in practical applications.
In recent years, transformer-based models [78, 57, 82] and Vision-Language Models (VLMs) [2, 12, 8], have achieved impressive success in HOI detection field. These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information. Despite these advancements, such models remain highly sensitive to various real-world corruptions, e.g., occlusion, lighting variations, and environmental shifts, which hinder their robustness and limit practical deployment. For instance, minor corruptions can cause substantial drops in accuracy, which poses significant risks in safety-critical applications like autonomous driving or robotics. This sensitivity raises serious concerns about the robustness and reliability of these models in unpredictable environments. Ensuring robust and reliable HOI detection in diverse, complex corruptions remains an unsolved challenge.
Existing HOI detection methods [78, 57, 2] predominantly evaluate model performance in idealized, clean conditions, overlooking real-world robustness.
This gap in the evaluation pipeline highlights the urgent need for a dedicated robustness benchmark. Such a benchmark can systematically evaluate model performance under diverse corruptions and perturbations, offering key insights into resilience and adaptability.
In this paper, we introduce Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse corruptions designed to simulate real-world conditions based on HICO-DET [3] and V-COCO [13] datasets and robustness-focused metrics including Mean Robustness Index (MRI) [18] and our novel Composite Robustness Index (CRI). RoHOI extends robustness evaluation beyond clean datasets, exposing model vulnerabilities and guiding future improvements. Through our evaluations of existing HOI detection methods on the newly proposed robustness test sets, we observe that their performance remains limited. This finding underscores the necessity of developing more robust HOI detection models to ensure greater reliability in real-world applications.
To mitigate the issue of performance degradation, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) method, which effectively leverages both holistic and partial cues from the clean training set. Specifically, we utilize SAM [28] to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying ratios. This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process. Our proposed learning strategy delivers superior performance in HOI detection, achieving state-of-the-art (SOTA) performance on clean test set while delivering substantial improvements in recognition accuracy under the evaluation with the proposed corruptions. This advancement highlights the model’s enhanced resilience and reliability in challenging real-world conditions for HOI detection. The main contributions of this paper are summarized as follows:
  • We for the first time introduce a robustness benchmark specifically tailored for HOI detection. Our benchmark incorporates $20$ distinct corruptions that closely mirror real-world challenges and a set of robustness-focused metrics.
  • We propose a novel semantic-aware masking curriculum learning approach that adaptively guides the model to optimize learning on both holistic and partial information.
  • SAMPL yields SOTA in terms of robustness, surpassing prior HOI models under both clean and corrupted conditions. Our benchmark establishes a comprehensive evaluation framework for robustness in HOI detection.
2 Related Work
2.1 Human Object Interaction Detection
Human-object interaction (HOI) detection [67, 11, 14, 32, 59, 47, 65, 69, 77, 81] seeks to automatically identify and interpret relationships between humans and objects in images. Existing approaches can be clustered into two major groups, which are one-stage and two-stage approaches.
Two-Stage HOI Detection. Two-stage methods [51, 64, 34, 80] rely on the detection results from the existing object detectors, e.g., Faster R-CNN [52]. CNN-based HOI detection methods [64, 80, 51] are explored by the researchers in the early stage. Graph structures [80, 64, 58, 61, 73] are well explored in the past to construct knowledge relation reasoning to pursue better human object interaction detection. Wan et al. [59] proposed pose aware multi-level network for HOI task. Zhang et al. [74] proposed an unpaired transformer architecture to achieve a better capture of the human-object interaction relations. Li et al. [33] proposed integrating and decomposing networks for HOI detection.
One-stage HOI Detection. One-stage HOI detection approaches [70, 75] have gained traction due to their efficiency and end-to-end learning advantages over two-stage methods. Liao et al. [35] introduced PPDM, leveraging parallel point detection and matching for real-time HOI detection. Kim et al. [24] proposed UnionDet, introducing a union-level detector to directly predict interaction triplets. Zhong et al. [79] introduced a glance-and-gaze mechanism to adaptively model action-aware points and refine interaction predictions. Recent advancements in transformer-based architectures further improved one-stage HOI detection. Kim et al. [27] proposed MUREN, which incorporates relational context learning to refine interaction reasoning. Yuan et al. [70] introduced RLIPv2, integrating Asymmetric Language-Image Fusion (ALIF) for scalable relational pretraining, achieving SOTA performance. VLMs have also been explored to enhance HOI detection, improving semantic understanding and generalization across unseen interactions. Liu et al. [23] proposed VLM-HOI, leveraging large-scale vision-language models for improved generalization. Zhang et al. [30] introduced EZ-HOI, enhancing zero-shot HOI triplet prediction through VLM-based feature learning. Addressing robustness in HOI detection, Wang et al. [7] proposed UAHOI, integrating uncertainty estimation to improve stability under varying conditions.
Most existing works are evaluated on clean datasets without considering data corruptions. However, in real-world applications, HOI detection models encounter various challenges leading to significant performance degradation. To address this, we introduce the first robustness benchmark for HOI detection, incorporating $20$ test-time corruptions.
2.2 Visual Robustness Benchmarks
Robustness is crucial for visual understanding as it ensures models can maintain high performance despite variations in environmental conditions, such as noise, occlusions, or lighting changes, which are common in real-world scenarios [4, 60, 53, 54, 1, 39, 68, 22, 44, 15, 63, 16, 40, 46, 71]. Researchers have constructed diverse robustness benchmarks to assess the capability of models coming from different visual understanding fields under diverse perturbations. Liu et al. [38] proposed ARES-Bench benchmark to evaluate the robustness of existing image classification methods under perturbations. Ishma et al. [21] constructed a large-scale benchmark for the visual question answer task. Hendrycks et al. [18] proposed ImageNet-C, a widely used benchmark assessing model robustness against common corruptions like noise, blur, and weather effects. Li et al. [31] proposed ImageNet-E to benchmark the model robustness in attribute editing field. Peng et al. [49] proposed new benchmark to asses robustness of existing one-shot skeleton based human action recognition models against occlusion perturbation. Chen et al. [9] constructed RoDLA to enable the evaluation of robust document layout analysis. However, none of the existing works focus on the robust evaluation of HOI detection methods. In this work, we open the vistas of the uncharted robust HOI detection area by constructing new RoHOI benchmark on two public HOI detection datasets and put forward the SAMPL method to achieve more reliable HOI detection performances.
3 Benchmark
3.1 Corruptions
To rigorously assess HOI detection robustness, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under structured real-world corruptions. While many corruptions in existing datasets [18], such as brightness and contrast shifts, are mitigated through standard augmentation, their weather-based effects (e.g., synthetic snow and frost) often fail to capture real-world optical complexities. Our benchmark introduces 20 corruption types, categorized into four groups: Optical system-induced (OS) artifacts, Sensor, compression, and transmission (SCT) artifacts , Environmentally induced (EI) artifacts, and Geometric and scene (G&S) distortions (see Fig. 1). Each corruption is applied at five severity levels, enabling a fine-grained analysis of robustness under increasing degradation.
OS Artifacts degrade spatial clarity and obscure interaction details by distorting object boundaries. Our benchmark includes motion blur (MB), defocus blur (DB), and Gaussian blur (GauB), which simulate degradation caused by camera movement, misfocus, and signal processing. We refine glass blur (GB) to better replicate distortions caused by imaging through uneven or scratched glass surfaces.
SCT Artifacts stem from hardware limitations, lossy encoding, and data transmission errors. Alongside standard corruptions like Gaussian noise (GauN), shot noise (ShN), salt-and-pepper noise (S&P), JPEG artifacts (JPEG), and speckle noise (SN), we introduce packet loss (PL), which simulates data degradation due to incomplete transmissions.
EI Artifacts alter image quality due to external conditions. Extreme lighting variations cause exposure (EXP) distortion. The rainbow effect (RE) misaligns color channels, distorting object hues. Occlusion (OCC) removes critical visual information when foreign objects block key scene elements, while vignette effect (VE) creates uneven brightness, reducing visibility at image borders.
G&S Distortions alter spatial relationships and object proportions. Moire patterns (MP) create artificial wave-like distortions in high-frequency textures. Screen crack (SC) introduce irregular fractures and glare artifacts, while elastic transformation (ET) warp object structures. Perspective distortion (PD) skews object shapes when captured at extreme angles. Pixelation (PIX) replaces fine details with block-like artifacts due to low-resolution rendering, and zoom blur (ZB) degrades object sharpness through rapid motion along the optical axis.
3.2 Datasets
To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO. HICO-DET [62] contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories. Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios. V-COCO [13] is a subset of MS-COCO [37] with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes. Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions. Evaluating these datasets under corruption provides a systematic robustness assessment beyond clean-image performance.
3.3 Baselines
To systematically evaluate the robustness of HOI detection models, we benchmark both two-stage and one-stage architectures, including transformer-based and vision-language pretraining approaches. Two-stage methods first detect objects before predicting interactions. We evaluate QPIC [56] and UPT [74], both of which use object detectors to extract human-object proposals before interaction classification. We include CDN [72], which refines interaction reasoning via a cascade structure, and DiffHOI [66], which leverages text-to-image diffusion priors to improve rare interaction recognition. One-stage methods directly predict interactions without explicit object detection, improving efficiency. We evaluate FGAHOI [43], which enhances spatial alignment using fine-grained anchors, and QAHOI [5], which integrates query-based attention for interaction localization. Additionally, we include MUREN [27], which refines relational context modeling, and SOV-STG [6], which optimizes decoder attention for verb-object association. GEN-VLKT [36] incorporates visual-language pretraining to improve interaction recognition. RLIPv2 [70] incorporates region-language alignment for fine-grained interaction reasoning. Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning. Our benchmark systematically evaluates these models across diverse corruption types and severity levels, providing a comprehensive assessment of HOI detection robustness.
3.4 Evaluation Metrics
We evaluate clean test set of HICO-DET [62] using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare ($<$10 instances), and non-rare ($\geq$10 instances) categories [70]. For clean test set of V-COCO [13], we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$. To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI). MRI, following the definition in prior work [18], measures average performance under corruptions but overlooks stability across severity levels. To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
3.4.1 Mean Robustness Index (MRI)
MRI serves as a baseline measure of a model’s average performance across corruption types and severities:
$\text{MRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{1}{L_{c}}\sum_{l=1}^{L_{c}}mAP_{c,l}\right),$ (1)
where $C$ denotes the number of corruption types, $L_{c}$ represents the number of severity levels for the $c$-th corruption, and $mAP_{c,l}$ is the model’s performance at level $l$ of corruption $c$. While MRI provides an overall robustness estimate, it does not account for performance fluctuations across corruption levels.
3.4.2 Composite Robustness Index (CRI)
To address MRI’s limitations, we propose CRI, which penalizes performance instability across corruption levels while normalizing against clean-set performance:
$\text{CRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{\overline{mAP}_{c}}{mAP_{clean}}\cdot\frac{1}{\log(1+\sigma_{c})+1}\right),$ (2)
where $\overline{mAP}c$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels. The clean-set performance $mAP{clean}$ is defined as mAP (Full) for HICO-DET and $\text{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5. The penalty term $\frac{1}{\log(1+\sigma_{c})+1}$ reduces scores for models with high variance, ensuring that robustness accounts for both overall performance and stability.
By integrating absolute accuracy with degradation consistency, CRI favors models that maintain stable performance across corruption levels, mitigating the impact of abrupt performance drops. For completeness, we also report standard HICO-DET mAP and V-COCO role-based AP scores, following prior benchmarks (detailed in Supplementary).
4 Method
In this section, we present our novel approach, Semantic-Aware Masking-based Progressive Learning (SAMPL), which improves model optimization by enhancing both holistic and partial feature learning. This is achieved through dynamically masked inputs, with the mask generation guided by the SAM [28] to identify Regions Of Interest (ROI) in HOI detection. The adaptive progressive learning strategy adjusts the masking procedure automatically based on the model’s performance on the validation set and the recorded frequency, ensuring more robust feature learning. We will now provide a detailed explanation of the design of our approach.
SAM Guided HOI-Aware Semantic Masking (SAMSM). Our goal is to maximize the utilization of the information provided by the clean training set to enhance the model’s robustness. This is achieved by ensuring that both holistic and partial cues are effectively extracted and learned, ultimately benefiting model optimization. By applying masks of varying scales to the semantic regions of interest in the image (i.e., areas related to interacted objects and person areas), we assist the model in learning HOI cues within the restricted regions.
To implement this masking strategy, we use ground-truth bounding boxes and SAM-generated masks for instance-aware occlusions. Each object’s bounding box is processed by SAM to obtain a fine-grained segmentation mask $\mathbf{M}_{b}$. To prevent reliance on instance contours, we apply random edge expansion, ensuring occlusion does not strictly follow the instance’s boundaries. Instead of naively occluding entire bounding boxes, we selectively mask regions within $\mathbf{M}_{b}$ while preserving surrounding contextual cues. This strategy prevents the model from relying solely on object appearance and encourages it to infer interactions based on partial information. During training, the masked pixels in the image are replaced with zero values.
Next, we define various severity levels for the mask to obtain different partial cues from the original image. These severity levels are represented by $\Omega=\{\omega_{1},\omega_{2},\omega_{3},\omega_{4}\}$, corresponding to clean, low, middle, and high levels. The clean level represents the original, unmasked data to support holistic feature learning. For the other three levels, cover ratios are determined to control the final height and width of the generated black box for semantic-aware masking, which can be denoted as $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$, where $\omega\in\left[\omega_{2},\omega_{3},\omega_{4}\right]$.
We then generate the mask for the severity level $\omega$ using the following equation:
$\mathbf{M}^{b}_{\omega}=\Phi(ConvHull(Dilation(\mathbf{M}_{b})),\left[r^{w}_{\omega},r^{h}_{\omega}\right]),$ (3)
In this process, we first apply binary dilation to the semantic mask $\mathbf{M}_{b}$, then compute the rectangular convex hull, and finally resize the resulting rectangular mask using the corresponding ratio $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$. The operation $\Phi$ refers to the resizing procedure that controls the mask size. The final mask, $\mathbf{M}^{b}_{\omega}$, is the resulting mask that aids partial feature learning in our proposed robust training method.
Algorithm 1 SAMPL
Input: Severity levels of semantic aware masking: $\Omega=\{\omega_{1},\omega_{2},\omega_{3},\omega_{4}\}$, initial threshold: $\tau_{\text{init}}$, maximum epochs: $T$, HOI detection model: $\mathcal{H}_{\theta}$.
Initialize: $N(\Omega)\leftarrow\mathbf{1}^{4\times 1}$, $\Delta Q_{\max}\leftarrow-\infty$, $p\leftarrow 2$.
1:for $t=1$ to $T$ do
2:     Compute $Q_{\omega}(t)$ for $\omega\in\{\omega_{1},\omega_{p}\}$ using $\mathcal{H}_{\theta}$.
3:     Get $S_{\omega_{p}}(t)=\sum_{i}{N({\omega_{i}})}\cdot Q_{\omega_{p}}(t)$. Get $S_{\omega_{1}}(t)=N({\omega_{1}})\cdot Q_{\omega_{1}}(t)$.
4:     Select severity level $\omega^{*}=\arg\min_{\omega\in\{\omega_{1},\omega_{p}\}}S_{\omega}(t)$.
5:     if $\omega^{*}\neq\omega_{1}$ then
6:         Compute $\Delta Q(t)=\frac{Q_{\omega_{p}}(t)-Q_{\omega_{p}}(t-1)}{Q_{\omega_{p}}(t-1)}$.
7:         Update $\Delta Q_{\max}=\max(\Delta Q_{\max},|\Delta Q(t)|)$.
8:         Compute $\tau(t)=\tau_{\text{init}}\cdot\frac{\Delta Q_{\max}}{|\Delta Q(t)|+\epsilon}$.
9:         if $|\Delta Q(t)|<\tau(t)$ and $p<4$ then
10:              $p:=p+1$. Train model on $\omega_{p}$.
11:              $N(\omega_{p}):=N(\omega_{p})+1$. $\Delta Q_{\max}\leftarrow-\infty$
12:         else  Train model on $\omega^{*}$. $N(\omega_{p}):=N(\omega_{p})+1$.
13:         end if
14:     else  Train model on $\omega_{1}$. $N(\omega_{1}):=N(\omega_{1})+1$.
15:     end if
16:end for
[…]
Two-Stage methods
QPIC [56] R50 29.07 21.85 31.23 13.13 0.21 58.80 60.98 30.04 0.19
QPIC [56] R101 29.90 23.92 31.69 14.45 0.22 58.27 60.74 32.05 0.19
UPT [74] R50 31.66 25.94 33.36 17.77 0.23 58.96 64.47 35.92 0.15
UPT [74] R101 32.31 28.55 33.44 17.44 0.24 60.70 66.20 33.80 0.14
CDN-B [72] R50 31.78 27.55 33.05 15.41 0.22 62.29 64.42 34.30 0.19
CDN-L [72] R101 32.07 27.19 33.53 17.05 0.24 63.91 65.89 37.65 0.20
DiffHOI [66] R50 34.41 31.07 35.40 21.09 0.23 - - - -
One-Stage methods
FGAHOI [43] Swin-T 29.81 22.17 32.09 16.05 0.22 - - - -
GEN-VLKT-S [36] R50 33.75 29.25 35.10 17.01 0.22 62.41 64.46 34.69 0.19
GEN-VLKT-L [36] R101 34.95 31.18 36.08 18.93 0.23 63.58 65.93 37.45 0.20
QAHOI [5] Swin-T 28.47 22.44 30.27 15.07 0.25 58.23 58.70 29.33 0.20
SOV-STG-S [6] R50 33.80 29.28 35.15 16.94 0.22 62.88 64.47 34.40 0.20
MUREN [27] R50 32.88 28.70 34.13 16.32 0.23 68.10 69.88 37.10 0.18
RLIPv2-ParSeDA [70] Swin-T 38.60 33.66 40.07 24.58 0.28 68.83 70.76 45.05 0.24
SAMPL (ours) Swin-T 39.16 33.89 40.73 24.25 0.29 69.04 71.13 48.83 0.27
[…]
Score Guided Progressive Feature Learning. To fully exploit the potential of the proposed SAM-guided semantic masks for maximizing HOI cue extraction, we introduce a novel score-guided progressive feature learning framework that incorporates a frequency memory bank across all severity levels.
Initially, training begins with the lowest severity level, while a memory bank is designed to track the utilization frequency of different severity levels. The perturbation level is then progressively increased to a more challenging level based on the relative change in the evaluation metric between consecutive epochs. Notably, the severity level is only upgraded when the selected level differs from $\omega_{1}$. Let $\Delta Q(t)$ represent the relative change in the evaluation score on the validation set at epoch $t$:
$\Delta Q(t)=\frac{Q_{\omega}(t)-Q_{\omega}(t-1)}{Q_{\omega}(t-1)},$ (4)
where $Q_{\omega}(t-1)$ is the evaluation metric for the current masking severity level in the previous epoch. To ensure stability, the perturbation level is upgraded when the magnitude of $\Delta Q(t)$ falls below a dynamically adjusted threshold $\tau(t)$:
$\tau(t)=\tau_{\text{init}}\cdot\frac{\Delta Q_{\max}}{|\Delta Q(t)|+\epsilon},$ (5)
where $\tau_{\text{init}}$ is the initial threshold, $\Delta Q_{\max}=\max_{k\leq t}|\Delta Q(k)|$ is the maximum observed relative change in $Q$ up to epoch $t$, and $\epsilon$ is a small constant to avoid division by zero.
The severity level $l$ is updated as:
$p:=\begin{cases}p+1&\text{if }|\Delta Q(t)|<\tau(t)\text{ and}~{}p<4,\\ p&\text{otherwise}.\end{cases}$ (6)
At epoch $t$, let $N(\omega)$ represent the number of times one severity level $\omega$ has been selected up to epoch $t$ to serve as the aforementioned frequency memory bank. Let $Q_{\omega}(t)$ denote the validation performance for severity level $\omega$ at epoch $t$. The severity level score $S_{\omega}(t)$ is thereby defined as:
$S_{\omega_{p}}(t)=\sum_{i}{N({\omega_{i}})}\cdot Q_{\omega_{p}}(t),\omega_{i}\in\Omega/\{\omega_{1}\}.$ (7)
Then we obtain $S_{\omega_{1}}=N({\omega_{1}})\cdot Q_{\omega_{1}}(t)$. At each epoch, the next training severity level is chosen between the original data $\omega_{1}$ (i.e., lowest severity level to achieve holistic feature learning) and $\omega_{p}$. The selected severity level $\omega^{*}$ is determined as:
$\omega^{*}=\arg\min_{\omega\in\{\omega_{1},\omega_{p}\}}S_{\omega}(t).$ (8)
The whole procedure is summarized in Alg. 1.
5 Experimental Results
5.1 Implementation Details
[…]
5.2 Analysis of the Benchmark
Figure 2: Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks. The left bar chart groups performance by corruption categories, while the right radar chart compares models across five severity levels for three corruptions. Corruptions are denoted as Abbreviation, i.e., $L\gamma$, where $\gamma$ represents the severity level. Details are shown in Sec. 3.1.
Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET [62] and V-COCO [13] datasets. SAMPL surpasses RLIPv2 [70], improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations. On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy. These results are further supported by Fig. 2, where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the supplementary). While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods. These results highlight SAMPL’s effectiveness in enhancing HOI detection under corrupted conditions through our proposed semantic-aware masking and progressive learning. Through a comprehensive analysis of representative HOI models, we derive the following key findings.
Observation 1: Large-scale vision-language pretraining significantly enhances the robustness of transformer-based models. According to the experiments conducted both on the clean test set and our perturbed test set on the two datasets, shown in Table 1, we observe that transformer-based architectures exhibit varying levels of robustness depending on their pretraining strategies. RLIPv2 [70], which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT [36], SOV-STG [6] and other transformer-based methods. This suggests that large-scale vision-language pretraining provides a more structured relational feature space, mitigating feature drift and localization sensitivity in HOI detection tasks. Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT [74], show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 [70]. The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images. Missed predictions are in dashed red, false interactions in red, and correct ones in blue.
Observation 2: Object localization strategies dictate robustness in two-stage architectures. Two-stage methods show varying robustness across different object localization strategies. Region-proposal-based models [72] achieve high clean-set accuracy but degrade under corruption due to error propagation in localization. In contrast, query-based methods [74] adapt object representations dynamically, enhancing robustness on HICO-DET. However, this advantage diminishes on V-COCO, where UPT with a larger backbone exhibits lower CRI than CDN [72] with a smaller backbone, likely due to V-COCO’s limited interaction diversity. This suggests that query-based detection benefits from diverse datasets but struggles with constrained interaction distributions.
[…]
AP#1_role AP#2_role
RLIPv2-ParSeDA$+$SAMSM 68.27 70.24 48.20 0.27
RLIPv2-ParSeDA$+$AGNP 68.33 70.51 44.74 0.24
RLIPv2-ParSeDA$+$RPP 67.69 69.64 47.59 0.25
Table 2: Ablation study on various perturbation strategies during training. We evaluate SAM-guided HOI-aware Semantic Masking (SAMSM), Adaptive Gaussian Noise Perturbation (AGNP), and Randomized Pixelation Perturbation (RPP) on the V-COCO.
Observation 3: Interaction representation entanglement amplifies robustness degradation under perturbations. Models with tightly coupled interaction representations achieve higher clean-set accuracy but are more prone to corruption due to feature distortion propagation. MUREN [27], with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness. Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s [5] adaptive query-based approach improves robustness by dynamically refining predictions. This underscores a key trade-off: stronger feature coupling boosts accuracy but reduces resilience to corruptions.
Observation 4: Dataset interaction diversity directly influences robustness generalization. Models trained on HIO-DET achieve higher MRI and CRI than V-COCO-trained models due to greater interaction diversity, enhancing resilience. However, SAMPL shows stronger robustness gains on V-COCO, as its progressive learning strategy compensates for dataset limitations, achieving comparable robustness with less data and fewer iterations. This underscores the importance of diverse interactions for general robustness and adaptive training for faster improvements in smaller datasets. Structural design considerations in both architectures and training paradigms are crucial for balancing recognition accuracy and robustness in real-world applications.
[…]
AP#1_role AP#2_role
40%/50%/60% 69.04 71.13 48.83 0.27
40%/60%/70% 68.51 70.53 48.44 0.27
50%/60%/80% 68.21 70.29 48.26 0.27
50%/70%/80% 68.52 70.45 45.72 0.22
Table 3: Ablation study on training with different severity level combinations. SAMSM masks each instance based on 40%–80% of its bounding box height and width, using three candidate levels during training.
5.3 Analysis of the Ablation Study
To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset. The results, presented in Tables 2 and 3, analyze the effects of different perturbation types and perturbation level settings on model performance.
Ablation of the Semantic-Aware Masking: Table 2 evaluates three perturbation strategies applied independently during training, excluding our progressive learning method to isolate their direct impact. SAM-guided HOI aware Semantic Masking (SAMSM) is applied in 50% of training epochs, occluding instance’s regions while preserving contextual cues. Adaptive Gaussian Noise Perturbation (AGNP) introduces feature-level noise with a 50% probability, while Randomized Pixelation Perturbation (RPP) pixelates input images with the same probability as data augmentation. SAMSM achieves the highest MRI and CRI scores, suggesting that structured occlusions enhance robustness by enforcing relational reasoning under missing object regions. Unlike AGNP, which perturbs latent features without spatial structure, SAMSM provides localized challenges that align with real-world occlusions. Despite expectations that RPP would improve robustness—given that pixelation causes the most significant performance drop in baseline corruption tests (as shown in Fig. 4)—its randomized nature indiscriminately distorts both essential and redundant information, making adaptation less effective. These results suggest that robustness in HOI detection benefits more from structured perturbations that challenge object-awareness rather than indiscriminate degradations that introduce confounding artifacts.
Figure 4: Performance degradation of RLIPv2 [70] on HICO-DET [62] and V-COCO [13] under RoHOI. The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO. Clean dataset performance serves as a reference, with shaded areas indicating corruption impact, with abbreviations mentioned in Sec. 3.1.
Effect of Perturbation Levels: Table 3 analyzes the impact of masking severity on robustness. SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores. Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($>70\%$) disrupts spatial structures, reducing discriminative capacity. These findings emphasize the need for controlled occlusions that challenge models without excessive information loss.
Overall, perturbation-based training enhances HOI detection robustness, with SAMSM proving most effective by balancing global context preservation and local feature resilience. Moderate perturbation achieves optimal robustness without sacrificing clean-data accuracy.
5.4 Analysis of the Qualitative Results
Fig. 3 compares SAMPL with RLIPv2 [70] under various corruptions, demonstrating SAMPL’s superior robustness in the first four cases while highlighting shared failure patterns in the last two. Under Screen Crack corruption (Column 1), SAMPL accurately detects occluded background persons and their “ride” interactions, whereas RLIPv2 misses smaller subjects, indicating SAMPL’s stronger resilience to fragmented textures. In Pixelation noise (Column 2), our method correctly recognizes “work on computer”, while RLIPv2 misclassifies it as “cut”, suggesting greater robustness against structured perturbations. Under Salt-and-Pepper Noise (Column 3), RLIPv2 confuses “sit” with “ride”, likely due to the presence of an adjacent animal, whereas SAMPL maintains stable interaction recognition, reducing reliance on object-centric biases. In Moire Pattern (Column 4) corruption, RLIPv2 incorrectly detects “talk on phone”, possibly due to moire-induced distortions interfering with high-frequency texture recognition. However, under high-severity Glass Blur (Column $5$) and Moire Pattern corruption (Column 6), both models misclassify actions as “surf” or “talk on phone”, potentially because V-COCO’s limited action categories encourage overfitting to common interactions under extreme degradation rather than a traditional long-tail distribution issue. These results highlight SAMPL’s advantages in structured occlusions and texture-based corruptions. However, both models remain vulnerable under severe degradations, where dataset priors dominate predictions.
6 Conclusion
In this paper, we introduced RoHOI, the first robustness benchmark for HOI detection, evaluating models under $20$ real-world perturbations. Our analysis shows that existing models suffer significant degradation, highlighting the need for more robust approaches. To address this, we propose SAMPL, a Semantic-Aware Masking-based Progressive Learning method that utilizes SAM-generated region-aware masks and dynamically adjusts masking severity. SAMPL enhances robustness while maintaining high accuracy on clean datasets, outperforming state-of-the-art methods. By bridging the gap between clean-data performance and real-world robustness, our work advances reliable and adaptable HOI detection models for practical use.
Appendix A Social Impact and Limitations
Social Impact: We introduce RoHOI, the first benchmark for assessing HOI detection robustness under real-world corruptions, incorporating 20 degradation types to bridge the gap between theoretical advancements and deployment. Robust HOI detection is crucial for safety-critical applications, yet existing models struggle with occlusions and noise. To address this, we propose SAMPL, a semantic-aware masking and progressive learning approach that enhances resilience to distortions, outperforming SOTA models in both clean and corrupted settings. However, challenges remain, including biased predictions and disparities across interaction categories, underscoring the need for further research in bias mitigation, domain adaptation, and fairness-aware robustness.
Limitations: RoHOI focuses on synthetic corruptions but may not fully capture real-world domain shifts like scene-specific occlusions and adversarial attacks, highlighting the need for real-world corrupted data. While SAMPL improves robustness, it still struggles with severe distortions (e.g., moire patterns, glass blur), indicating the need for further architectural advancements.
Appendix B Standard HOI Evaluation Metrics
This section provides a detailed explanation of the evaluation metrics used for HICO-DET [64] and V-COCO [13] in our main paper. These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
[…]
We evaluate models on HICO-DET using the standard mAP under an IoU threshold of 0.5. The reported metrics include mAP (Full), which evaluates all 600 HOI categories; mAP (Rare), assessing performance on low-frequency categories with fewer than 10 instances; and mAP (Non-Rare), focusing on categories with at least 10 instances.
B.2 $\mathrm{AP}^{\#1}_{\text{role}}$ and $\mathrm{AP}^{\#2}_{\text{role}}$
We evaluate V-COCO using $\mathrm{AP}_{\text{role}}$, which measures interaction detection accuracy. $\mathrm{AP}^{\#1}_{\text{role}}$ assesses action recognition without strict role object localization. $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5.
Model MB DB GauB GB GauN ShN S&P JPEG SN PL EXP RE OCC VE MP SC ET PD PIX ZB
CDN-L [72] 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
GEN-VLKT-L [36] 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
MUREN [27] 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
RLIPv2-ParSeDA [70] 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
SAMPL (Ours) 26.63 19.03 58.56 13.87 52.46 68.06 52.66 65.81 42.24 46.69 69.22 57.31 62.04 70.77 47.89 60.72 48.99 53.21 29.90 30.49
Table 4: Comparison of MRI scores between our SAMPL method and the best-performing one-stage and two-stage HOI detection models on the RoHOI benchmark. The table presents results across 20 corruption types, evaluating model robustness under diverse corruptions. Corruption abbreviations follow the definitions provided in the main paper.
Level 1 Level 2 Level 3 Level 4 Level 5
Motion Blur (MB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Defocus Blur (DB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gaussian Blur (GauB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Glass Blur (GB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 5: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
Appendix C Corruption Types in the Dataset
Building upon the general introduction in the main text, this section provides a more detailed breakdown of the $20$ corruption types used in our RoHOI benchmark.
C.1 Optical System-Induced (OS) Artifacts
Motion Blur (MB): Blur caused by the relative motion between the camera and the scene during exposure. This artifact is included as it commonly occurs in dynamic scenes or when handheld cameras are used. Severity levels are generated by increasing the kernel size of a linear motion blur filter. (Details shown in Table 6, row $1$.)
Defocus Blur (DB): Blur resulting from the lens being improperly focused on the subject. This noise reflects the limitations of camera focus systems, particularly in low-light or high-speed scenarios. Severity levels are controlled by varying the radius of a circular point spread function (PSF). (Details shown in Fig. 6, row $2$.)
Gaussian Blur (GauB): Blur due to scattering or inherent optical limitations of the lens. This artifact simulates the diffusion of light caused by imperfect optical surfaces or aperture settings. Severity levels are adjusted by increasing the standard deviation of the Gaussian kernel. (Details shown in Fig. 6, row $3$.)
Glass Blur (GB): Blur introduced when imaging through imperfect or contaminated transparent materials (e.g., glass or plastic). This type is included to represent common real-world challenges such as photographing through windows or protective covers. Severity levels are generated by applying iterative small displacements and local blurring. (Details shown in Fig. 6, row $4$.)
Level 1 Level 2 Level 3 Level 4 Level 5
Salt-and-Pepper Noise (S&P) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
JPEG Artifact (JPEG) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Speckle Noise (SN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Gaussian Noise (GauN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Shot Noise (ShN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Packet Loss (PL) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 6: Visual examples of Sensor, Compression and Transmission (SCT) artifacts across severity levels.
C.2 Sensor, Compression and Transmission (SCT) Artifacts
Salt-and-Pepper Noise (S&P): Discrete black-and-white pixel noise caused by sensor malfunction or transmission errors. This represents rare but impactful noise patterns in hardware-limited imaging systems. Severity levels are generated by increasing the probability of random pixel corruption. (Details shown in Fig. 6, row $1$.)
JPEG Artifacts (JPEG): Blocky or blurry artifacts resulting from lossy image compression, which are used to evaluate model robustness to low-quality image formats frequently encountered in data storage and transmission. Severity levels are controlled by progressively reducing the JPEG quality factor. (Details shown in Fig. 6, row $2$.)
Speckle Noise (SN): Speckle noise resembles the grainy texture seen on old printed photos or low-quality digital images taken in dim lighting. At lower severity levels, it appears as faint specks of dust on a camera lens, barely noticeable unless viewed closely. As severity increases, the noise becomes more pronounced, resembling the rough texture of sandpaper or the graininess of a poorly printed newspaper. At the highest severity levels, the effect is akin to heavy rain blurring a car windshield, significantly degrading image clarity. (Details shown in Fig. 6, row $3$.)
Gaussian Noise (GauN): Additive noise caused by sensor imperfections such as thermal noise or quantization errors. This is a ubiquitous noise model, reflecting baseline sensor limitations in most imaging devices. Severity levels are determined by progressively increasing the noise standard deviation. (Details shown in Fig. 6, row $4$.)
Shot Noise (ShN): Random variations in pixel intensity caused by photon-counting statistics during exposure. This noise type represents the fundamental physical limits of image sensors in low-light conditions. Severity levels are controlled by adjusting the scale of the Poisson-distributed noise. (Details shown in Fig. 6, row $5$.)
Packet Loss (PL): Missing or distorted image regions caused by data corruption during transmission. This noise simulates real-world challenges in wireless or network-based imaging systems. Severity levels are generated by increasing the number and size of corrupted or duplicated regions. (Details shown in Fig. 6, row $6$.)
C.3 Environmentally Induced (EI) Artifacts
Level 1 Level 2 Level 3 Level 4 Level 5
Exposure (EXP) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Rainbow Effect (RE) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Occlusion (OCC) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Vignette Effect (VE) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 7: Visual examples of Environmentally Induced (EI) artifacts across severity levels.
Exposure Artifacts (EXP): Loss of image details caused by underexposure or overexposure due to extreme lighting conditions. Higher severity levels increase the intensity of overexposure (bright regions saturate) or underexposure (dark regions lose details), leading to stronger visual degradation. This type is included to represent real-world challenges in scenes with poor or uncontrolled lighting. (Details shown in Fig. 7, row $1$.)
Rainbow Effect (RE): Typically caused by the interaction between the projection system’s optical components (e.g., color wheels in DLP projectors) and the camera sensor. As severity increases, the color distortion becomes more pronounced, with broader and more intense rainbow bands appearing across the image, significantly altering local color distribution. This noise enables the evaluation of model robustness against chromatic distortions, particularly in scenes involving projected displays and reflective surfaces. (Details shown in Fig. 7, row $2$.)
Occlusion (OCC): Partial obstruction of the field of view by external objects. Increasing severity levels introduce randomly placed black rectangular occlusions of varying sizes, progressively removing more contextual information from the image. This is a critical inclusion as occlusions are prevalent in crowded or dynamic environments. (Details shown in Fig. 7, row $3$.)
Vignette Effect (VE): Gradual darkening at the edges of an image caused by lens design or aperture settings. With higher severity, the vignette effect has darker and wider peripheral regions, increasingly suppressing details in the outer parts of the image. This type represents a common optical limitation that affects peripheral details in wide-field imaging. (Details shown in Fig. 7, row $4$.)
C.4 Geometric and Scene (G&S) Distortions
Level 1 Level 2 Level 3 Level 4 Level 5
Moire Pattern (MP) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Screen Crack (SC) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Elastic Transform (ET) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Perspective Distortion (PD) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Pixelation (PIX) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Zoom Blur (ZB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 8: Visual examples of Geometric and Scene (G&S) Distortions across severity levels.
Moire Pattern (MP): Interference patterns resulting from fine texture overlap with the camera’s pixel grid (e.g., fabric or screens). This artifact captures the challenges posed by imaging repetitive textures or high-frequency patterns. Severity levels adjust the density and contrast of the interference patterns. (Details shown in Fig. 8, row $1$.)
Screen Crack Artifacts (SC): Visual obstructions or patterns caused by photographing through cracked or damaged surfaces. This type is included to reflect realistic scenarios where images are captured through broken screens or fractured glass. Higher severity levels introduce progressively denser and larger crack patterns, further obscuring the image. (Details shown in Fig. 8, row $2$.)
Elastic Transform (ET): Non-linear distortions caused by perspective changes, material deformations, or uneven air density. This artifact represents challenges encountered in scenes with flexible or moving objects. Higher severity levels correspond to greater displacement magnitudes, leading to noticeable warping. (Details shown in Fig. 8, row $3$.)
Perspective Distortion (PD): This artifact simulates geometric distortions from lens imperfections, such as barrel and pincushion distortion. Severity is increased by applying progressively stronger transformation matrices, exaggerating the curvature and altering the perceived structure of objects. (Details shown in Fig. 8, row $4$.)
Pixelation (PIX): Blocky appearance due to low resolution or excessive digital zoom. This represents a key limitation in devices with constrained sensor capabilities or in images subjected to aggressive downscaling. Higher severity levels result in coarser pixel blocks, significantly reducing image detail. (Details shown in Fig. 8, row $5$.)
Zoom Blur (ZB): Radial blur caused by rapid changes in focal length during image capture. Zoom blur is simulated by iteratively applying slight scaling transformations and averaging the results to create a progressive radial blur effect. Higher severity levels increase both the scaling intensity and the number of blended layers, leading to a stronger blur radiating outward from the center. This method ensures a smooth and natural-looking zoom effect without abrupt artifacts. (Details shown in Fig. 8, row $6$.)
C.5 Purpose of Inclusion
These noise types are selected to comprehensively represent real-world scenarios that affect image quality, ranging from optical and sensor limitations to environmental and dynamic scene factors. Their inclusion ensures the dataset’s relevance for evaluating model robustness in diverse and realistic conditions.
Appendix D Benchmark Results on Corruption Types
Table 4 presents the MRI scores of our SAMPL method compared to the strongest one-stage and two-stage HOI detection baselines on the RoHOI benchmark. SAMPL consistently outperforms prior methods across all 20 corruption types, achieving the highest scores in every category. Notably, our approach shows significant improvements in handling severe degradations such as pixelation and packet loss, demonstrating enhanced robustness. These results validate the effectiveness of our SAMPL strategy in improving model resilience against real-world corruptions.

### v2

@@
 To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges.
-Despite advances, current models struggle with environmental variability, occlusion, and noise.
-Our benchmark, RoHOI, includes 20 corruption types based on HICO-DET and V-COCO datasets and a new robustness-focused metric.
-We systematically analyze existing models in the related field, revealing significant performance drops under corruptions.
-To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, dynamically adjusting the model’s optimization to enhance robust feature learning.
-Extensive experiments show our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection.
+Despite advances, current models struggle with environmental variability, occlusions, and noise.
+Our benchmark, RoHOI, includes 20 corruption types based on the HICO-DET and V-COCO datasets and a new robustness-focused metric.
+We systematically analyze existing models in the HOI field, revealing significant performance drops under corruptions.
+To improve robustness, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) strategy to guide the model to be optimized based on holistic and partial cues, thus dynamically adjusting the model’s optimization to enhance robust feature learning.
+Extensive experiments show that our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection.
 Benchmarks, datasets, and code will be made publicly available at https://github.com/Kratos-Wen/RoHOI.
-1 Introduction
-Human-Object Interaction (HOI) detection is a key task in visual scene understanding, involving the identification and interpretation of interactions between humans and objects in an image [25, 19, 78, 56, 26, 20].
-It serves as a foundation for many critical applications, including autonomous driving [45, 48], robotics [65], video surveillance [10, 50], and augmented reality [76], where a fine-grained understanding of human behavior and object interactions is paramount [13].
-Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups, i.e., Optical system-induced artifacts, Sensor, compression and transmission artifacts, Environmentally induced artifacts, and Geometric and scene distortions.
+ [^1] The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation).
+This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026).
+Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups.
 Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions.
-These corruptions simulate real-world degradation factors encountered in practical applications.
-In recent years, transformer-based models [78, 57, 82] and Vision-Language Models (VLMs) [2, 12, 8], have achieved impressive success in HOI detection field.
+These corruptions simulate realistic semantic and structural disturbances specifically encountered in practical application.
+Introduction
+Human-Object Interaction (HOI) detection aims to identify and interpret interactions between humans and objects (Kim et al.
+2021; Hou et al.
+2021a; Zhang et al.
+2022; Tamura, Ohashi, and Yoshinaga 2021; Kim et al.
+2022; Hou et al.
+2021b).
+It serves many critical applications, including autonomous driving (Martin et al.
+2019; Peng et al.
+2022), robotics (Xu et al.
+2020), video surveillance (Dogariu et al.
+2020; Prest, Ferrari, and Schmid 2013), and augmented reality (Zhang, Cui, and Fang 2025), where a fine-grained understanding of human behavior and object handling is paramount (Gupta and Malik 2015).
+In recent years, transformer-based models (Zhang et al.
+2022; Tu et al.
+2023; Zou et al.
+2021) and Vision-Language Models (VLMs) (Cao et al.
+2023; Gao et al.
+2024; Chen et al.
+2024a), have achieved impressive success in the HOI detection field.
 These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information.
-Despite these advancements, such models remain highly sensitive to various real-world corruptions, e.g., occlusion, lighting variations, and environmental shifts, which hinder their robustness and limit practical deployment.
-For instance, minor corruptions can cause substantial drops in accuracy, which poses significant risks in safety-critical applications like autonomous driving or robotics.
-This sensitivity raises serious concerns about the robustness and reliability of these models in unpredictable environments.
-Ensuring robust and reliable HOI detection in diverse, complex corruptions remains an unsolved challenge.
-Existing HOI detection methods [78, 57, 2] predominantly evaluate model performance in idealized, clean conditions, overlooking real-world robustness.
-This gap in the evaluation pipeline highlights the urgent need for a dedicated robustness benchmark.
-Such a benchmark can systematically evaluate model performance under diverse corruptions and perturbations, offering key insights into resilience and adaptability.
-In this paper, we introduce Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse corruptions designed to simulate real-world conditions based on HICO-DET [3] and V-COCO [13] datasets and robustness-focused metrics including Mean Robustness Index (MRI) [18] and our novel Composite Robustness Index (CRI).
-RoHOI extends robustness evaluation beyond clean datasets, exposing model vulnerabilities and guiding future improvements.
-Through our evaluations of existing HOI detection methods on the newly proposed robustness test sets, we observe that their performance remains limited.
-This finding underscores the necessity of developing more robust HOI detection models to ensure greater reliability in real-world applications.
-To mitigate the issue of performance degradation, we propose a Semantic-Aware Masking-based Progressive Learning (SAMPL) method, which effectively leverages both holistic and partial cues from the clean training set.
-Specifically, we utilize SAM [28] to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying ratios.
+Despite these advancements, existing models remain sensitive to real-life corruptions, particularly semantic distortions like occlusions and perspective shifts, as well as environmental variations.
+Even minor disturbances may pose significant risks in critical applications like autonomous driving and robotics.
+Enhancing robustness under these conditions remains a crucial but unresolved challenge.
+Existing HOI detection methods (Zhang et al.
+2022; Tu et al.
+2023; Cao et al.
+2023) had their performance evaluated predominantly under ideal conditions.
+Thus, a dedicated robustness benchmark tailored explicitly to HOI detection, covering both pixel-level and semantic-level corruptions, is urgently required to put their resilience to the test.
+In this paper, we introduce the Robustness benchmark for HOI detection (RoHOI), systematically assessing HOI detection models under $20$ diverse forms of data corruption.
+These were designed to simulate real-world conditions.
+The HICO-DET (Chao et al.
+2018) and V-COCO (Gupta and Malik 2015) datasets served as starting point.
+We include robustness-focused metrics, such as the Mean Robustness Index (MRI) (Hendrycks and Dietterich 2019) and our novel Composite Robustness Index (CRI).
+Through our evaluations of existing HOI detection methods on the newly proposed robustness test set, we observe that their performance is less than optimal.
+This finding underscores the need to develop more robust HOI detection models.
+To mitigate the issue of performance degradation, we propose our Semantic-Aware Masking-based Progressive Learning (SAMPL) method.
+Specifically, we utilize SAM (Kirillov et al.
+2023) to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying size ratios.
 This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process.
-Our proposed learning strategy delivers superior performance in HOI detection, achieving state-of-the-art (SOTA) performance on clean test set while delivering substantial improvements in recognition accuracy under the evaluation with the proposed corruptions.
-This advancement highlights the model’s enhanced resilience and reliability in challenging real-world conditions for HOI detection.
+Our proposed learning strategy delivers superior HOI detection, retaining state-of-the-art (SOTA) performance on a clean test set while delivering substantial improvements under corruptions.
+This highlights the model’s enhanced resilience and reliability in challenging real-world conditions.
 The main contributions of this paper are summarized as follows:
-  • We for the first time introduce a robustness benchmark specifically tailored for HOI detection.
+  • We for the first time introduce a robustness benchmark specifically tailored to HOI detection.
 Our benchmark incorporates $20$ distinct corruptions that closely mirror real-world challenges and a set of robustness-focused metrics.
@@
   • SAMPL yields SOTA in terms of robustness, surpassing prior HOI models under both clean and corrupted conditions.
-Our benchmark establishes a comprehensive evaluation framework for robustness in HOI detection.
-2 Related Work
-2.1 Human Object Interaction Detection
-Human-object interaction (HOI) detection [67, 11, 14, 32, 59, 47, 65, 69, 77, 81] seeks to automatically identify and interpret relationships between humans and objects in images.
-Existing approaches can be clustered into two major groups, which are one-stage and two-stage approaches.
-Two-Stage HOI Detection.
-Two-stage methods [51, 64, 34, 80] rely on the detection results from the existing object detectors, e.g., Faster R-CNN [52].
-CNN-based HOI detection methods [64, 80, 51] are explored by the researchers in the early stage.
-Graph structures [80, 64, 58, 61, 73] are well explored in the past to construct knowledge relation reasoning to pursue better human object interaction detection.
-Wan et al.
-[59] proposed pose aware multi-level network for HOI task.
-Zhang et al.
-[74] proposed an unpaired transformer architecture to achieve a better capture of the human-object interaction relations.
-Li et al.
-[33] proposed integrating and decomposing networks for HOI detection.
-One-stage HOI Detection.
-One-stage HOI detection approaches [70, 75] have gained traction due to their efficiency and end-to-end learning advantages over two-stage methods.
-Liao et al.
-[35] introduced PPDM, leveraging parallel point detection and matching for real-time HOI detection.
-Kim et al.
-[24] proposed UnionDet, introducing a union-level detector to directly predict interaction triplets.
-Zhong et al.
-[79] introduced a glance-and-gaze mechanism to adaptively model action-aware points and refine interaction predictions.
-Recent advancements in transformer-based architectures further improved one-stage HOI detection.
-Kim et al.
-[27] proposed MUREN, which incorporates relational context learning to refine interaction reasoning.
-Yuan et al.
-[70] introduced RLIPv2, integrating Asymmetric Language-Image Fusion (ALIF) for scalable relational pretraining, achieving SOTA performance.
-VLMs have also been explored to enhance HOI detection, improving semantic understanding and generalization across unseen interactions.
-Liu et al.
-[23] proposed VLM-HOI, leveraging large-scale vision-language models for improved generalization.
-Zhang et al.
-[30] introduced EZ-HOI, enhancing zero-shot HOI triplet prediction through VLM-based feature learning.
-Addressing robustness in HOI detection, Wang et al.
-[7] proposed UAHOI, integrating uncertainty estimation to improve stability under varying conditions.
-Most existing works are evaluated on clean datasets without considering data corruptions.
-However, in real-world applications, HOI detection models encounter various challenges leading to significant performance degradation.
-To address this, we introduce the first robustness benchmark for HOI detection, incorporating $20$ test-time corruptions.
-2.2 Visual Robustness Benchmarks
-Robustness is crucial for visual understanding as it ensures models can maintain high performance despite variations in environmental conditions, such as noise, occlusions, or lighting changes, which are common in real-world scenarios [4, 60, 53, 54, 1, 39, 68, 22, 44, 15, 63, 16, 40, 46, 71].
-Researchers have constructed diverse robustness benchmarks to assess the capability of models coming from different visual understanding fields under diverse perturbations.
-Liu et al.
-[38] proposed ARES-Bench benchmark to evaluate the robustness of existing image classification methods under perturbations.
-Ishma et al.
-[21] constructed a large-scale benchmark for the visual question answer task.
-Hendrycks et al.
-[18] proposed ImageNet-C, a widely used benchmark assessing model robustness against common corruptions like noise, blur, and weather effects.
-Li et al.
-[31] proposed ImageNet-E to benchmark the model robustness in attribute editing field.
-Peng et al.
-[49] proposed new benchmark to asses robustness of existing one-shot skeleton based human action recognition models against occlusion perturbation.
-Chen et al.
-[9] constructed RoDLA to enable the evaluation of robust document layout analysis.
-However, none of the existing works focus on the robust evaluation of HOI detection methods.
-In this work, we open the vistas of the uncharted robust HOI detection area by constructing new RoHOI benchmark on two public HOI detection datasets and put forward the SAMPL method to achieve more reliable HOI detection performances.
-3 Benchmark
-3.1 Corruptions
-To rigorously assess HOI detection robustness, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under structured real-world corruptions.
-While many corruptions in existing datasets [18], such as brightness and contrast shifts, are mitigated through standard augmentation, their weather-based effects (e.g., synthetic snow and frost) often fail to capture real-world optical complexities.
-Our benchmark introduces 20 corruption types, categorized into four groups: Optical system-induced (OS) artifacts, Sensor, compression, and transmission (SCT) artifacts , Environmentally induced (EI) artifacts, and Geometric and scene (G&S) distortions (see Fig.
-1).
-Each corruption is applied at five severity levels, enabling a fine-grained analysis of robustness under increasing degradation.
-OS Artifacts degrade spatial clarity and obscure interaction details by distorting object boundaries.
-Our benchmark includes motion blur (MB), defocus blur (DB), and Gaussian blur (GauB), which simulate degradation caused by camera movement, misfocus, and signal processing.
+Related Work
+Human Object Interaction Detection.
+Human-object interaction (HOI) detection (Yao and Fei-Fei 2010; Fang et al.
+2018; Gupta, Schwing, and Hoiem 2019; Li et al.
+2019; Wan et al.
+2019; Park et al.
+2018; Xu et al.
+2020; Yu, Liu, and Odobez 2018; Zhang et al.
+2015; Zhou et al.
+2020) aims to automatically identify and interpret relationships between humans and objects, typically divided into two categories: two-stage and one-stage approaches.
+Two-stage HOI detection methods (Qi et al.
+2018; Xu et al.
+2019; Liang et al.
+2021; Zhou and Chi 2019) first detect humans and objects independently before classifying interactions.
+Early CNN-based approaches (Xu et al.
+2019; Zhou and Chi 2019; Qi et al.
+2018) mainly utilized spatial contexts for interaction modeling.
+Later graph-based methods (Zhou and Chi 2019; Xu et al.
+2019; Ulutan, Iftekhar, and Manjunath 2020; Wang et al.
+2021; Zhang, Campbell, and Gould 2021) improved relational reasoning explicitly.
+Subsequent methods further enhanced precision via pose-aware representations (Wan et al.
+2019), transformer-based unpaired modeling (Zhang, Campbell, and Gould 2022), and integration-decomposition strategies (Li et al.
+2020).
+One-stage HOI detection methods (Yuan et al.
+2023; Zhang et al.
+2023) achieve efficiency by directly predicting human-object interactions end-to-end.
+Early methods exploited parallel point detection (Liao et al.
+2020) and union-level modeling (Kim et al.
+2020), while subsequent approaches refined interaction localization via adaptive glance-and-gaze mechanisms (Zhong et al.
+2021).
+Transformer-based methods enhanced relational context modeling (Kim, Jung, and Cho 2023) and language-image fusion (RLIPv2)(Yuan et al.
+2023), significantly improving performance.
+Recently, vision-language models (VLMs) were leveraged to boost generalization across unseen interactions (VLM-HOI(Kang et al.
+2024)) and zero-shot triplet predictions (EZ-HOI (Lei, Wang, and Tan 2024)).
+Robustness-specific methods further integrated uncertainty estimation (UAHOI (Chen, Chen, and Yang 2024)) to improve stability under perturbations.
+Nevertheless, all these approaches primarily evaluate performance on clean data, neglecting real-world corruptions.
+To address this gap, we introduce RoHOI, the first HOI robustness benchmark encompassing $20$ test-time corruptions.
+Visual Robustness Benchmarks.
+Robustness benchmarks have been widely established to assess visual model resilience across various tasks (Chattopadhyay et al.
+2021; Wang et al.
+2024; Sarkar et al.
+2024; Schiappa et al.
+2024; An et al.
+2024; Liu et al.
+2024b; Yi et al.
+2024; Kamann and Rother 2020; Ma et al.
+2024b; Hao et al.
+2024; Xie et al.
+2025; Hao et al.
+2025; Liu et al.
+2024c; Michaelis et al.
+2019; Zeng et al.
+2024), including image classification (ARES-Bench (Liu et al.
+2024a)), visual question answering (Ishmam et al.
+2024), common corruption evaluation (ImageNet-C (Hendrycks and Dietterich 2019)), attribute editing robustness (ImageNet-E (Li et al.
+2023)), one-shot skeleton-based action recognition under occlusion (Peng et al.
+2023), and robust document layout analysis (RoDLA (Chen et al.
+2024b)).
+However, robustness benchmarks explicitly designed for HOI detection have not been explored.
+In this work, we address the uncharted robust HOI detection area by constructing the new RoHOI benchmark and by putting forward the SAMPL method to achieve more reliable HOI detection performance.
+Benchmark
+Corruptions.
+To rigorously assess the robustness of HOI detection, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under real-world corruptions.
+Existing benchmarks such as ImageNet-C (Hendrycks and Dietterich 2019) primarily cover low-level pixel disturbances like brightness or blur.
+In contrast, RoHOI explicitly includes semantic-level corruptions specifically tailored to challenge spatial and relational reasoning critical for HOI detection.
+RoHOI explicitly introduces 20 corruptions derived from HICO-DET and V-COCO images, categorized into four groups: optical system artifacts such as motion blur and glass blur; sensor, compression, and transmission distortions such as packet loss and JPEG artifacts; environmental effects including exposure imbalance and occlusion; and geometric or structural disruptions such as screen cracks, elastic deformation, and perspective distortion.
+The corruptions range from pixel-level noise to semantic-level challenges that affect spatial alignment and interaction reasoning.
+Each corruption is implemented with five severity levels to support detailed analysis under progressively more difficult conditions.
+Optical System-induced (OS) Artifacts degrade spatial clarity and obscure interaction details by distorting object boundaries.
+Our benchmark includes motion blur (MB), defocus blur (DB), and Gaussian blur (GauB), which simulate degradation caused by camera movement, defocus, and signal processing.
 We refine glass blur (GB) to better replicate distortions caused by imaging through uneven or scratched glass surfaces.
-SCT Artifacts stem from hardware limitations, lossy encoding, and data transmission errors.
+Sensor, Compression and Transmission (SCT) Artifacts stem from hardware limitations, lossy encoding, and data transmission errors.
 Alongside standard corruptions like Gaussian noise (GauN), shot noise (ShN), salt-and-pepper noise (S&P), JPEG artifacts (JPEG), and speckle noise (SN), we introduce packet loss (PL), which simulates data degradation due to incomplete transmissions.
-EI Artifacts alter image quality due to external conditions.
+Environmentally Induced (EI) Artifacts alter image quality due to external conditions.
 Extreme lighting variations cause exposure (EXP) distortion.
 The rainbow effect (RE) misaligns color channels, distorting object hues.
-Occlusion (OCC) removes critical visual information when foreign objects block key scene elements, while vignette effect (VE) creates uneven brightness, reducing visibility at image borders.
-G&S Distortions alter spatial relationships and object proportions.
+Occlusion (OCC) removes critical visual information when foreign objects block key scene elements, while the vignette effect (VE) creates uneven brightness, reducing visibility at image borders.
+Geometric and Scene (G&S) Distortions alter spatial relationships and object proportions.
 Moire patterns (MP) create artificial wave-like distortions in high-frequency textures.
@@
 Pixelation (PIX) replaces fine details with block-like artifacts due to low-resolution rendering, and zoom blur (ZB) degrades object sharpness through rapid motion along the optical axis.
-3.2 Datasets
+Datasets.
 To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO.
-HICO-DET [62] contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories.
+HICO-DET (Wang et al.
+2020) contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories.
 Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios.
-V-COCO [13] is a subset of MS-COCO [37] with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes.
+V-COCO (Gupta and Malik 2015) is a subset of MS-COCO (Lin et al.
+2014) with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes.
 Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions.
 Evaluating these datasets under corruption provides a systematic robustness assessment beyond clean-image performance.
-3.3 Baselines
+Baselines.
 To systematically evaluate the robustness of HOI detection models, we benchmark both two-stage and one-stage architectures, including transformer-based and vision-language pretraining approaches.
 Two-stage methods first detect objects before predicting interactions.
-We evaluate QPIC [56] and UPT [74], both of which use object detectors to extract human-object proposals before interaction classification.
-We include CDN [72], which refines interaction reasoning via a cascade structure, and DiffHOI [66], which leverages text-to-image diffusion priors to improve rare interaction recognition.
+We evaluate QPIC (Tamura, Ohashi, and Yoshinaga 2021) and UPT (Zhang, Campbell, and Gould 2022), both of which use object detectors to extract human-object proposals before interaction classification.
+We include CDN (Zhang et al.
+2021), which refines interaction reasoning via a cascade structure, and DiffHOI (Yang et al.
+2023), which leverages text-to-image diffusion priors to improve rare interaction recognition.
 One-stage methods directly predict interactions without explicit object detection, improving efficiency.
-We evaluate FGAHOI [43], which enhances spatial alignment using fine-grained anchors, and QAHOI [5], which integrates query-based attention for interaction localization.
-Additionally, we include MUREN [27], which refines relational context modeling, and SOV-STG [6], which optimizes decoder attention for verb-object association.
-GEN-VLKT [36] incorporates visual-language pretraining to improve interaction recognition.
-RLIPv2 [70] incorporates region-language alignment for fine-grained interaction reasoning.
+We evaluate FGAHOI (Ma et al.
+2024a), which enhances spatial alignment using fine-grained anchors, and QAHOI (Chen and Yanai 2023), which integrates query-based attention for interaction localization.
+Additionally, we include MUREN (Kim, Jung, and Cho 2023), which refines relational context modeling, and SOV-STG (Chen, Wang, and Yanai 2023), which optimizes decoder attention for verb-object association.
+GEN-VLKT (Liao et al.
+2022) incorporates visual-language pretraining to improve interaction recognition.
+RLIPv2 (Yuan et al.
+2023) incorporates region-language alignment for fine-grained interaction reasoning.
 Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning.
 Our benchmark systematically evaluates these models across diverse corruption types and severity levels, providing a comprehensive assessment of HOI detection robustness.
-3.4 Evaluation Metrics
-We evaluate clean test set of HICO-DET [62] using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare ($<$10 instances), and non-rare ($\geq$10 instances) categories [70].
-For clean test set of V-COCO [13], we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$.
+Evaluation Metrics.
+We evaluate the clean test set of HICO-DET (Wang et al.
+2020) using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare (<<10 instances), and non-rare ($\geq$10 instances) categories (Yuan et al.
+2023).
+For the clean test set of V-COCO (Gupta and Malik 2015), we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$.
 To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI).
-MRI, following the definition in prior work [18], measures average performance under corruptions but overlooks stability across severity levels.
+MRI, following the definition in prior work (Hendrycks and Dietterich 2019), measures average performance under corruptions but overlooks stability across severity levels.
 To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
-3.4.1 Mean Robustness Index (MRI)
-MRI serves as a baseline measure of a model’s average performance across corruption types and severities:
+Mean Robustness Index (MRI) serves as a baseline measure of a model’s average performance across corruption types and severities:
 $\text{MRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{1}{L_{c}}\sum_{l=1}^{L_{c}}mAP_{c,l}\right),$ (1)
@@
 While MRI provides an overall robustness estimate, it does not account for performance fluctuations across corruption levels.
-3.4.2 Composite Robustness Index (CRI)
-To address MRI’s limitations, we propose CRI, which penalizes performance instability across corruption levels while normalizing against clean-set performance:
+Composite Robustness Index (CRI) addresses MRI’s limitations. It penalizes performance instability across corruption levels while normalizing against clean-set performance:
 $\text{CRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{\overline{mAP}_{c}}{mAP_{clean}}\cdot\frac{1}{\log(1+\sigma_{c})+1}\right),$ (2)
-where $\overline{mAP}c$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels.
-The clean-set performance $mAP{clean}$ is defined as mAP (Full) for HICO-DET and $\text{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5.
+where $\overline{mAP}_{c}$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels.
+The clean-set performance $mAP_{clean}$ is defined as mAP (Full) for HICO-DET and $\mathrm{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5.
 The penalty term $\frac{1}{\log(1+\sigma_{c})+1}$ reduces scores for models with high variance, ensuring that robustness accounts for both overall performance and stability.
@@
 For completeness, we also report standard HICO-DET mAP and V-COCO role-based AP scores, following prior benchmarks (detailed in Supplementary).
-4 Method
-In this section, we present our novel approach, Semantic-Aware Masking-based Progressive Learning (SAMPL), which improves model optimization by enhancing both holistic and partial feature learning.
-This is achieved through dynamically masked inputs, with the mask generation guided by the SAM [28] to identify Regions Of Interest (ROI) in HOI detection.
-The adaptive progressive learning strategy adjusts the masking procedure automatically based on the model’s performance on the validation set and the recorded frequency, ensuring more robust feature learning.
-We will now provide a detailed explanation of the design of our approach.
+Method
+In this section, we present our novel Semantic-Aware Masking-based Progressive Learning (SAMPL), aiming to enhance model robustness by progressively guiding models from holistic feature reliance towards robust contextual inference.
+Unlike standard augmentation methods, SAMPL employs adaptive semantic masking guided by SAM (Kirillov et al.
+2023), dynamically adjusting masking severity based on validation performance.
+This structured semantic-level masking strategy explicitly promotes robust contextual reasoning by training models to infer interactions from progressively restricted semantic regions, thus effectively improving general robustness against a variety of realistic disturbances.
+Moreover, SAMPL’s structured progressive curriculum inherently regularizes feature representation, effectively reducing overfitting to dataset biases, thereby further improving generalization under unforeseen corruptions.
 SAM Guided HOI-Aware Semantic Masking (SAMSM).
-Our goal is to maximize the utilization of the information provided by the clean training set to enhance the model’s robustness.
-This is achieved by ensuring that both holistic and partial cues are effectively extracted and learned, ultimately benefiting model optimization.
-By applying masks of varying scales to the semantic regions of interest in the image (i.e., areas related to interacted objects and person areas), we assist the model in learning HOI cues within the restricted regions.
-To implement this masking strategy, we use ground-truth bounding boxes and SAM-generated masks for instance-aware occlusions.
-Each object’s bounding box is processed by SAM to obtain a fine-grained segmentation mask $\mathbf{M}_{b}$.
-To prevent reliance on instance contours, we apply random edge expansion, ensuring occlusion does not strictly follow the instance’s boundaries.
-Instead of naively occluding entire bounding boxes, we selectively mask regions within $\mathbf{M}_{b}$ while preserving surrounding contextual cues.
-This strategy prevents the model from relying solely on object appearance and encourages it to infer interactions based on partial information.
-During training, the masked pixels in the image are replaced with zero values.
-Next, we define various severity levels for the mask to obtain different partial cues from the original image.
-These severity levels are represented by $\Omega=\{\omega_{1},\omega_{2},\omega_{3},\omega_{4}\}$, corresponding to clean, low, middle, and high levels.
-The clean level represents the original, unmasked data to support holistic feature learning.
-For the other three levels, cover ratios are determined to control the final height and width of the generated black box for semantic-aware masking, which can be denoted as $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$, where $\omega\in\left[\omega_{2},\omega_{3},\omega_{4}\right]$.
-We then generate the mask for the severity level $\omega$ using the following equation:
+To fully leverage clean training data, we generate instance-aware masks based on ground-truth bounding boxes and SAM-derived fine-grained masks ($\mathbf{M}_{b}$).
+Random edge expansions are applied to prevent strict reliance on instance contours, selectively masking semantic regions while preserving contextual cues.
+During training, masked pixels are set to zero, compelling the model to infer interactions from partial visual information.
+We define various severity levels for the mask, represented by $\Omega=\{\omega_{1},\omega_{2},\omega_{3},\omega_{4}\}$, corresponding to clean, low, middle, and high levels.
+Clean level $\omega_{1}$ employs unmasked data for holistic feature learning, while other levels apply increasing occlusion severity via controlled cover ratios $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$.
+Each semantic mask $\mathbf{M}^{b}_{\omega}$ is computed as:
 $\mathbf{M}^{b}_{\omega}=\Phi(ConvHull(Dilation(\mathbf{M}_{b})),\left[r^{w}_{\omega},r^{h}_{\omega}\right]),$ (3)
-In this process, we first apply binary dilation to the semantic mask $\mathbf{M}_{b}$, then compute the rectangular convex hull, and finally resize the resulting rectangular mask using the corresponding ratio $\left[r^{w}_{\omega},r^{h}_{\omega}\right]$.
-The operation $\Phi$ refers to the resizing procedure that controls the mask size.
-The final mask, $\mathbf{M}^{b}_{\omega}$, is the resulting mask that aids partial feature learning in our proposed robust training method.
+where dilation and convex hull operations refine the mask shape, and $\Phi$ denotes resizing.
+This structured masking approach not only simulates realistic semantic occlusions but also inherently promotes context-aware reasoning, significantly enhancing robustness against various semantic and structural corruptions.
 Algorithm 1 SAMPL
@@
 Two-Stage methods
-QPIC [56] R50 29.07 21.85 31.23 13.13 0.21 58.80 60.98 30.04 0.19
-QPIC [56] R101 29.90 23.92 31.69 14.45 0.22 58.27 60.74 32.05 0.19
-UPT [74] R50 31.66 25.94 33.36 17.77 0.23 58.96 64.47 35.92 0.15
-UPT [74] R101 32.31 28.55 33.44 17.44 0.24 60.70 66.20 33.80 0.14
-CDN-B [72] R50 31.78 27.55 33.05 15.41 0.22 62.29 64.42 34.30 0.19
-CDN-L [72] R101 32.07 27.19 33.53 17.05 0.24 63.91 65.89 37.65 0.20
-DiffHOI [66] R50 34.41 31.07 35.40 21.09 0.23 - - - -
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 One-Stage methods
-FGAHOI [43] Swin-T 29.81 22.17 32.09 16.05 0.22 - - - -
-GEN-VLKT-S [36] R50 33.75 29.25 35.10 17.01 0.22 62.41 64.46 34.69 0.19
-GEN-VLKT-L [36] R101 34.95 31.18 36.08 18.93 0.23 63.58 65.93 37.45 0.20
-QAHOI [5] Swin-T 28.47 22.44 30.27 15.07 0.25 58.23 58.70 29.33 0.20
-SOV-STG-S [6] R50 33.80 29.28 35.15 16.94 0.22 62.88 64.47 34.40 0.20
-MUREN [27] R50 32.88 28.70 34.13 16.32 0.23 68.10 69.88 37.10 0.18
-RLIPv2-ParSeDA [70] Swin-T 38.60 33.66 40.07 24.58 0.28 68.83 70.76 45.05 0.24
-SAMPL (ours) Swin-T 39.16 33.89 40.73 24.25 0.29 69.04 71.13 48.83 0.27
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 […]
 Score Guided Progressive Feature Learning.
-To fully exploit the potential of the proposed SAM-guided semantic masks for maximizing HOI cue extraction, we introduce a novel score-guided progressive feature learning framework that incorporates a frequency memory bank across all severity levels.
+To effectively utilize SAM-guided masks, we design a novel score-guided progressive learning framework that incorporates a frequency memory bank across severity levels.
 Initially, training begins with the lowest severity level, while a memory bank is designed to track the utilization frequency of different severity levels.
-The perturbation level is then progressively increased to a more challenging level based on the relative change in the evaluation metric between consecutive epochs.
+The perturbation level is then progressively increased based on the relative change in the evaluation metric between consecutive epochs.
 Notably, the severity level is only upgraded when the selected level differs from $\omega_{1}$.
@@
 The whole procedure is summarized in Alg. 1.
-5 Experimental Results
-5.1 Implementation Details
-[…]
-5.2 Analysis of the Benchmark
+Experimental Results
+Implementation Details.
+Our SAMPL method is built upon RLIPv2 (Yuan et al.
+2023) with a Swin-Tiny (Liu et al.
+2021) backbone and has $214M$ trainable parameters.
+The backbone is pretrained on Visual Genome (Krishna et al.
+2017), MS-COCO (Lin et al.
+2014) and Object365 (Shao et al.
+2019), and fine-tuned on the HICO-DET or V-COCO datasets.
+We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of 1ee-4 (backbone and text encoder at 1ee-5), linear warmup followed by cosine decay, weight decay of 1ee-4, and batch size $2$ per GPU with gradient accumulation over $2$ steps.
+Models are trained for $80$ epochs using cross-entropy loss for object classification, focal loss for verb classification, and L1 + GIoU loss for bounding, with KL divergence for latent space regularization.
+The encoder and decoder consist of $6$ and $3$ layers, respectively.
+For the dynamic severity adjustment mechanism, we set the initial threshold as $\tau_{\text{init}}=0.15$ and $\epsilon=10^{-6}$.
 Figure 2: Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks.
@@
 Details are shown in Sec.
-3.1.
-Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET [62] and V-COCO [13] datasets.
-SAMPL surpasses RLIPv2 [70], improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations.
+Benchmark.
+Analysis of the Benchmark.
+Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET (Wang et al.
+2020) and V-COCO (Gupta and Malik 2015) datasets.
+SAMPL surpasses RLIPv2 (Yuan et al.
+2023), improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations.
 On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy.
@@
 2, where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the supplementary).
-While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods.
+While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods by progressively training the model with semantic-aware occlusions, enabling it to learn context-driven features that are resilient to diverse and severe corruptions.
 These results highlight SAMPL’s effectiveness in enhancing HOI detection under corrupted conditions through our proposed semantic-aware masking and progressive learning.
@@
 Observation 1: Large-scale vision-language pretraining significantly enhances the robustness of transformer-based models.
-According to the experiments conducted both on the clean test set and our perturbed test set on the two datasets, shown in Table 1, we observe that transformer-based architectures exhibit varying levels of robustness depending on their pretraining strategies.
-RLIPv2 [70], which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT [36], SOV-STG [6] and other transformer-based methods.
+Experiments on the clean and corrupted test set (Table 1) reveal that transformer-based architectures exhibit robustness variations dependent on their pretraining strategies.
+RLIPv2, which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT, SOV-STG and other transformer-based methods.
 This suggests that large-scale vision-language pretraining provides a more structured relational feature space, mitigating feature drift and localization sensitivity in HOI detection tasks.
-Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT [74], show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
-Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 [70].
+Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT, show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
+Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 (Yuan et al.
+2023).
 The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images.
@@
 Two-stage methods show varying robustness across different object localization strategies.
-Region-proposal-based models [72] achieve high clean-set accuracy but degrade under corruption due to error propagation in localization.
-In contrast, query-based methods [74] adapt object representations dynamically, enhancing robustness on HICO-DET.
-However, this advantage diminishes on V-COCO, where UPT with a larger backbone exhibits lower CRI than CDN [72] with a smaller backbone, likely due to V-COCO’s limited interaction diversity.
+Region-proposal-based models (Zhang et al.
+2021) achieve high clean-set accuracy but degrade under corruption due to error propagation in localization.
+In contrast, query-based methods (Zhang, Campbell, and Gould 2022) adapt object representations dynamically, enhancing robustness on HICO-DET.
+However, on V-COCO, UPT with a larger backbone surprisingly shows lower CRI than CDN with a smaller backbone, likely attributed to V-COCO’s limited interaction diversity.
 This suggests that query-based detection benefits from diverse datasets but struggles with constrained interaction distributions.
@@
 AP#1_role AP#2_role
-RLIPv2-ParSeDA$+$SAMSM 68.27 70.24 48.20 0.27
-RLIPv2-ParSeDA$+$AGNP 68.33 70.51 44.74 0.24
-RLIPv2-ParSeDA$+$RPP 67.69 69.64 47.59 0.25
+[…]
+[…]
+[…]
 Table 2: Ablation study on various perturbation strategies during training.
@@
 Models with tightly coupled interaction representations achieve higher clean-set accuracy but are more prone to corruption due to feature distortion propagation.
-MUREN [27], with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness.
-Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s [5] adaptive query-based approach improves robustness by dynamically refining predictions.
+MUREN, with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness.
+Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s adaptive query-based approach improves robustness by dynamically refining predictions.
 This underscores a key trade-off: stronger feature coupling boosts accuracy but reduces resilience to corruptions.
@@
 AP#1_role AP#2_role
-40%/50%/60% 69.04 71.13 48.83 0.27
-40%/60%/70% 68.51 70.53 48.44 0.27
-50%/60%/80% 68.21 70.29 48.26 0.27
-50%/70%/80% 68.52 70.45 45.72 0.22
+[…]
+[…]
+[…]
+[…]
 Table 3: Ablation study on training with different severity level combinations.
 SAMSM masks each instance based on 40%–80% of its bounding box height and width, using three candidate levels during training.
-5.3 Analysis of the Ablation Study
+Analysis of the Ablation Study.
 To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset.
@@
 These results suggest that robustness in HOI detection benefits more from structured perturbations that challenge object-awareness rather than indiscriminate degradations that introduce confounding artifacts.
-Figure 4: Performance degradation of RLIPv2 [70] on HICO-DET [62] and V-COCO [13] under RoHOI.
+Figure 4: Performance degradation of RLIPv2 (Yuan et al.
+2023) on HICO-DET and V-COCO under RoHOI.
 The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO.
-Clean dataset performance serves as a reference, with shaded areas indicating corruption impact, with abbreviations mentioned in Sec.
-3.1.
+Clean dataset performance serves as a reference, with shaded areas indicating corruption impact.
 Effect of Perturbation Levels: Table 3 analyzes the impact of masking severity on robustness.
 SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores.
-Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($>70\%$) disrupts spatial structures, reducing discriminative capacity.
+Moderate occlusions enhance feature learning via contextual cues, while excessive masking ($70\%$) disrupts spatial structures, reducing discriminative capacity.
 These findings emphasize the need for controlled occlusions that challenge models without excessive information loss.
@@
 Moderate perturbation achieves optimal robustness without sacrificing clean-data accuracy.
-5.4 Analysis of the Qualitative Results
+Analysis of the Qualitative Results.
 Fig.
-3 compares SAMPL with RLIPv2 [70] under various corruptions, demonstrating SAMPL’s superior robustness in the first four cases while highlighting shared failure patterns in the last two.
-Under Screen Crack corruption (Column 1), SAMPL accurately detects occluded background persons and their “ride” interactions, whereas RLIPv2 misses smaller subjects, indicating SAMPL’s stronger resilience to fragmented textures.
-In Pixelation noise (Column 2), our method correctly recognizes “work on computer”, while RLIPv2 misclassifies it as “cut”, suggesting greater robustness against structured perturbations.
-Under Salt-and-Pepper Noise (Column 3), RLIPv2 confuses “sit” with “ride”, likely due to the presence of an adjacent animal, whereas SAMPL maintains stable interaction recognition, reducing reliance on object-centric biases.
-In Moire Pattern (Column 4) corruption, RLIPv2 incorrectly detects “talk on phone”, possibly due to moire-induced distortions interfering with high-frequency texture recognition.
-However, under high-severity Glass Blur (Column $5$) and Moire Pattern corruption (Column 6), both models misclassify actions as “surf” or “talk on phone”, potentially because V-COCO’s limited action categories encourage overfitting to common interactions under extreme degradation rather than a traditional long-tail distribution issue.
+3 compares SAMPL with RLIPv2 under various corruptions, demonstrating SAMPL’s superior robustness in the first four cases while highlighting shared failure patterns in the last two.
+Under Screen Crack corruption (Col.
+$1$), SAMPL accurately detects occluded background persons and their “ride” interactions, whereas RLIPv2 misses smaller subjects, indicating SAMPL’s stronger resilience to fragmented textures, as SAMPL uses semantic-aware masking that selectively occludes HOI-relevant regions, training the model to recognize interactions from partial and degraded visual cues.
+In Pixelation noise (Col.
+$2$), our method correctly recognizes “work on computer”, while RLIPv2 misclassifies it as “cut”, suggesting greater robustness against structured perturbations.
+Under Salt-and-Pepper Noise (Col.
+$3$), RLIPv2 confuses “sit” with “ride”, likely due to the presence of an adjacent animal, whereas SAMPL maintains stable interaction recognition, reducing reliance on object-centric biases.
+In Moire Pattern (Col.
+$4$) corruption, RLIPv2 incorrectly detects “talk on phone”, possibly due to moire-induced distortions interfering with high-frequency texture recognition.
+However, under high-severity Glass Blur (Col.
+$5$) and Moire Pattern corruption (Col.
+$6$), both models misclassify actions as “surf” or “talk on phone”, potentially because V-COCO’s limited action categories encourage overfitting to common interactions under extreme degradation rather than a traditional long-tail distribution issue.
 These results highlight SAMPL’s advantages in structured occlusions and texture-based corruptions.
 However, both models remain vulnerable under severe degradations, where dataset priors dominate predictions.
-6 Conclusion
+Conclusion
 In this paper, we introduced RoHOI, the first robustness benchmark for HOI detection, evaluating models under $20$ real-world perturbations.
@@
 Appendix B Standard HOI Evaluation Metrics
-This section provides a detailed explanation of the evaluation metrics used for HICO-DET [64] and V-COCO [13] in our main paper.
+This section provides a detailed explanation of the evaluation metrics used for HICO-DET (Xu et al.
+2019) and V-COCO (Gupta and Malik 2015) in our main paper.
 These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
-[…]
+$\bm{\mathrm{mAP}}\,\textbf{(Full, Rare, Non-Rare)}$
 We evaluate models on HICO-DET using the standard mAP under an IoU threshold of 0.5.
 The reported metrics include mAP (Full), which evaluates all 600 HOI categories; mAP (Rare), assessing performance on low-frequency categories with fewer than 10 instances; and mAP (Non-Rare), focusing on categories with at least 10 instances.
-B.2 $\mathrm{AP}^{\#1}_{\text{role}}$ and $\mathrm{AP}^{\#2}_{\text{role}}$
+$\bm{\mathrm{AP}^{\#1}_{\text{role}}}$ and $\bm{\mathrm{AP}^{\#2}_{\text{role}}}$
 We evaluate V-COCO using $\mathrm{AP}_{\text{role}}$, which measures interaction detection accuracy.
@@
 Model MB DB GauB GB GauN ShN S&P JPEG SN PL EXP RE OCC VE MP SC ET PD PIX ZB
-CDN-L [72] 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
-GEN-VLKT-L [36] 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
-MUREN [27] 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
-RLIPv2-ParSeDA [70] 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
-SAMPL (Ours) 26.63 19.03 58.56 13.87 52.46 68.06 52.66 65.81 42.24 46.69 69.22 57.31 62.04 70.77 47.89 60.72 48.99 53.21 29.90 30.49
+CDN-L (Zhang et al. 2021) 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
+GEN-VLKT-L (Liao et al. 2022) 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
+MUREN (Kim, Jung, and Cho 2023) 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
+RLIPv2-ParSeDA (Yuan et al. 2023) 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
+[…]
 Table 4: Comparison of MRI scores between our SAMPL method and the best-performing one-stage and two-stage HOI detection models on the RoHOI benchmark.
@@
 Level 1 Level 2 Level 3 Level 4 Level 5
-Motion Blur (MB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Defocus Blur (DB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Gaussian Blur (GauB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Glass Blur (GB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 5: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
+Motion Blur (MB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Defocus Blur (DB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Gaussian Blur (GauB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Glass Blur (GB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 5: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
 Appendix C Corruption Types in the Dataset
 Building upon the general introduction in the main text, this section provides a more detailed breakdown of the $20$ corruption types used in our RoHOI benchmark.
-C.1 Optical System-Induced (OS) Artifacts
+Optical System-Induced (OS) Artifacts
 Motion Blur (MB): Blur caused by the relative motion between the camera and the scene during exposure.
@@
 Severity levels are generated by increasing the kernel size of a linear motion blur filter.
-(Details shown in Table 6, row $1$.)
+(Details shown in Figure 5, row $1$.)
 Defocus Blur (DB): Blur resulting from the lens being improperly focused on the subject.
@@
 Severity levels are controlled by varying the radius of a circular point spread function (PSF).
-(Details shown in Fig.
-6, row $2$.)
+(Details shown in Figure 5, row $2$.)
 Gaussian Blur (GauB): Blur due to scattering or inherent optical limitations of the lens.
@@
 Severity levels are adjusted by increasing the standard deviation of the Gaussian kernel.
-(Details shown in Fig.
-6, row $3$.)
+(Details shown in Figure 5, row $3$.)
 Glass Blur (GB): Blur introduced when imaging through imperfect or contaminated transparent materials (e.g., glass or plastic).
@@
 Severity levels are generated by applying iterative small displacements and local blurring.
-(Details shown in Fig.
-6, row $4$.)
+(Details shown in Figure 5, row $4$.)
 Level 1 Level 2 Level 3 Level 4 Level 5
-Salt-and-Pepper Noise (S&P) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-JPEG Artifact (JPEG) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Speckle Noise (SN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Gaussian Noise (GauN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Shot Noise (ShN) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Packet Loss (PL) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 6: Visual examples of Sensor, Compression and Transmission (SCT) artifacts across severity levels.
-C.2 Sensor, Compression and Transmission (SCT) Artifacts
+Salt-and-Pepper Noise (S&P) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+JPEG Artifact (JPEG) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Speckle Noise (SN) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Gaussian Noise (GauN) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Shot Noise (ShN) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Packet Loss (PL) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 6: Visual examples of Sensor, Compression and Transmission (SCT) artifacts across severity levels.
+Sensor, Compression and Transmission (SCT) Artifacts
 Salt-and-Pepper Noise (S&P): Discrete black-and-white pixel noise caused by sensor malfunction or transmission errors.
@@
 Severity levels are generated by increasing the probability of random pixel corruption.
-(Details shown in Fig.
-6, row $1$.)
+(Details shown in Figure 6, row $1$.)
 JPEG Artifacts (JPEG): Blocky or blurry artifacts resulting from lossy image compression, which are used to evaluate model robustness to low-quality image formats frequently encountered in data storage and transmission.
 Severity levels are controlled by progressively reducing the JPEG quality factor.
-(Details shown in Fig.
-6, row $2$.)
+(Details shown in Figure 6, row $2$.)
 Speckle Noise (SN): Speckle noise resembles the grainy texture seen on old printed photos or low-quality digital images taken in dim lighting.
@@
 At the highest severity levels, the effect is akin to heavy rain blurring a car windshield, significantly degrading image clarity.
-(Details shown in Fig.
-6, row $3$.)
+(Details shown in Figure 6, row $3$.)
 Gaussian Noise (GauN): Additive noise caused by sensor imperfections such as thermal noise or quantization errors.
@@
 Severity levels are determined by progressively increasing the noise standard deviation.
-(Details shown in Fig.
-6, row $4$.)
+(Details shown in Figure 6, row $4$.)
 Shot Noise (ShN): Random variations in pixel intensity caused by photon-counting statistics during exposure.
@@
 Severity levels are controlled by adjusting the scale of the Poisson-distributed noise.
-(Details shown in Fig.
-6, row $5$.)
+(Details shown in Figure 6, row $5$.)
 Packet Loss (PL): Missing or distorted image regions caused by data corruption during transmission.
@@
 Severity levels are generated by increasing the number and size of corrupted or duplicated regions.
-(Details shown in Fig.
-6, row $6$.)
-C.3 Environmentally Induced (EI) Artifacts
+(Details shown in Figure 6, row $6$.)
+Environmentally Induced (EI) Artifacts
 Level 1 Level 2 Level 3 Level 4 Level 5
-Exposure (EXP) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Rainbow Effect (RE) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Occlusion (OCC) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Vignette Effect (VE) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 7: Visual examples of Environmentally Induced (EI) artifacts across severity levels.
+Exposure (EXP) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Rainbow Effect (RE) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Occlusion (OCC) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Vignette Effect (VE) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 7: Visual examples of Environmentally Induced (EI) artifacts across severity levels.
 Exposure Artifacts (EXP): Loss of image details caused by underexposure or overexposure due to extreme lighting conditions.
@@
 This type is included to represent real-world challenges in scenes with poor or uncontrolled lighting.
-(Details shown in Fig.
-7, row $1$.)
+(Details shown in Figure 7, row $1$.)
 Rainbow Effect (RE): Typically caused by the interaction between the projection system’s optical components (e.g., color wheels in DLP projectors) and the camera sensor.
@@
 This noise enables the evaluation of model robustness against chromatic distortions, particularly in scenes involving projected displays and reflective surfaces.
-(Details shown in Fig.
-7, row $2$.)
+(Details shown in Figure 7, row $2$.)
 Occlusion (OCC): Partial obstruction of the field of view by external objects.
@@
 This is a critical inclusion as occlusions are prevalent in crowded or dynamic environments.
-(Details shown in Fig.
-7, row $3$.)
+(Details shown in Figure 7, row $3$.)
 Vignette Effect (VE): Gradual darkening at the edges of an image caused by lens design or aperture settings.
@@
 This type represents a common optical limitation that affects peripheral details in wide-field imaging.
-(Details shown in Fig.
-7, row $4$.)
-C.4 Geometric and Scene (G&S) Distortions
+(Details shown in Figure 7, row $4$.)
+Geometric and Scene (G&S) Distortions
 Level 1 Level 2 Level 3 Level 4 Level 5
-Moire Pattern (MP) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Screen Crack (SC) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Elastic Transform (ET) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Perspective Distortion (PD) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Pixelation (PIX) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Zoom Blur (ZB) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 8: Visual examples of Geometric and Scene (G&S) Distortions across severity levels.
+Moire Pattern (MP) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Screen Crack (SC) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Elastic Transform (ET) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Perspective Distortion (PD) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Pixelation (PIX) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Zoom Blur (ZB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 8: Visual examples of Geometric and Scene (G&S) Distortions across severity levels.
 Moire Pattern (MP): Interference patterns resulting from fine texture overlap with the camera’s pixel grid (e.g., fabric or screens).
@@
 Severity levels adjust the density and contrast of the interference patterns.
-(Details shown in Fig.
-8, row $1$.)
+(Details shown in Figure 8, row $1$.)
 Screen Crack Artifacts (SC): Visual obstructions or patterns caused by photographing through cracked or damaged surfaces.
@@
 Higher severity levels introduce progressively denser and larger crack patterns, further obscuring the image.
-(Details shown in Fig.
-8, row $2$.)
+(Details shown in Figure 8, row $2$.)
 Elastic Transform (ET): Non-linear distortions caused by perspective changes, material deformations, or uneven air density.
@@
 Higher severity levels correspond to greater displacement magnitudes, leading to noticeable warping.
-(Details shown in Fig.
-8, row $3$.)
+(Details shown in Figure 8, row $3$.)
 Perspective Distortion (PD): This artifact simulates geometric distortions from lens imperfections, such as barrel and pincushion distortion.
 Severity is increased by applying progressively stronger transformation matrices, exaggerating the curvature and altering the perceived structure of objects.
-(Details shown in Fig.
-8, row $4$.)
+(Details shown in Figure 8, row $4$.)
 Pixelation (PIX): Blocky appearance due to low resolution or excessive digital zoom.
@@
 Higher severity levels result in coarser pixel blocks, significantly reducing image detail.
-(Details shown in Fig.
-8, row $5$.)
+(Details shown in Figure 8, row $5$.)
 Zoom Blur (ZB): Radial blur caused by rapid changes in focal length during image capture.
@@
 This method ensures a smooth and natural-looking zoom effect without abrupt artifacts.
-(Details shown in Fig.
-8, row $6$.)
-C.5 Purpose of Inclusion
+(Details shown in Figure 8, row $6$.)
+Purpose of Inclusion
 These noise types are selected to comprehensively represent real-world scenarios that affect image quality, ranging from optical and sensor limitations to environmental and dynamic scene factors.

### v3

@@
 Human-Object Interaction (HOI) detection is crucial for robot-human assistance, enabling context-aware support.
-However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate prediction.
+However, models trained on clean datasets degrade in real-world conditions due to unforeseen corruptions, leading to inaccurate predictions.
 To address this, we introduce the first robustness benchmark for HOI detection, evaluating model resilience under diverse challenges.
@@
 Extensive experiments show that our approach outperforms state-of-the-art methods, setting a new standard for robust HOI detection.
-Benchmarks, datasets, and code will be made publicly available at https://github.com/Kratos-Wen/RoHOI.
- [^1] The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation).
-This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026).
-Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups.
-Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions.
-These corruptions simulate realistic semantic and structural disturbances specifically encountered in practical application.
-Introduction
-Human-Object Interaction (HOI) detection aims to identify and interpret interactions between humans and objects (Kim et al.
-2021; Hou et al.
-2021a; Zhang et al.
-2022; Tamura, Ohashi, and Yoshinaga 2021; Kim et al.
-2022; Hou et al.
-2021b).
-It serves many critical applications, including autonomous driving (Martin et al.
-2019; Peng et al.
-2022), robotics (Xu et al.
-2020), video surveillance (Dogariu et al.
-2020; Prest, Ferrari, and Schmid 2013), and augmented reality (Zhang, Cui, and Fang 2025), where a fine-grained understanding of human behavior and object handling is paramount (Gupta and Malik 2015).
-In recent years, transformer-based models (Zhang et al.
-2022; Tu et al.
-2023; Zou et al.
-2021) and Vision-Language Models (VLMs) (Cao et al.
-2023; Gao et al.
-2024; Chen et al.
-2024a), have achieved impressive success in the HOI detection field.
+Benchmarks, datasets, and code are available at https://github.com/KratosWen/RoHOI.
+1 Introduction
+Human-Object Interaction (HOI) detection aims to identify and interpret interactions between humans and objects Kim et al.
+(2021); Hou et al.
+(2021a); Zhang et al.
+(2022b); Tamura et al.
+(2021); Kim et al.
+(2022); Hou et al.
+(2021b).
+It serves many critical applications, including autonomous driving Martin et al.
+(2019); Peng et al.
+(2022), robotics Xu et al.
+(2020), video surveillance Dogariu et al.
+(2020); Prest et al.
+(2013), and augmented reality Zhang et al.
+(2025), where a fine-grained understanding of human behavior and object handling is paramount Gupta & Malik (2015).
+In recent years, transformer-based models Zhang et al.
+(2022b); Tu et al.
+(2023); Zou et al.
+(2021) and Vision-Language Models (VLMs) Cao et al.
+(2023); Gao et al.
+(2024); Chen et al.
+(2024b), have achieved impressive success in the HOI detection field.
 These models excel at capturing intricate relationships between human-object pairs by leveraging large-scale datasets and contextual information.
@@
 Enhancing robustness under these conditions remains a crucial but unresolved challenge.
-Existing HOI detection methods (Zhang et al.
-2022; Tu et al.
-2023; Cao et al.
-2023) had their performance evaluated predominantly under ideal conditions.
+Existing HOI detection methods Zhang et al.
+(2022b); Tu et al.
+(2023); Cao et al.
+(2023) had their performance evaluated predominantly under ideal conditions.
 Thus, a dedicated robustness benchmark tailored explicitly to HOI detection, covering both pixel-level and semantic-level corruptions, is urgently required to put their resilience to the test.
@@
 These were designed to simulate real-world conditions.
-The HICO-DET (Chao et al.
-2018) and V-COCO (Gupta and Malik 2015) datasets served as starting point.
-We include robustness-focused metrics, such as the Mean Robustness Index (MRI) (Hendrycks and Dietterich 2019) and our novel Composite Robustness Index (CRI).
+The HICO-DET Chao et al.
+(2018) and V-COCO Gupta & Malik (2015) datasets served as the starting point.
+We include robustness-focused metrics, such as the Mean Robustness Index (MRI) Hendrycks & Dietterich (2019) and our novel Composite Robustness Index (CRI).
 Through our evaluations of existing HOI detection methods on the newly proposed robustness test set, we observe that their performance is less than optimal.
@@
 To mitigate the issue of performance degradation, we propose our Semantic-Aware Masking-based Progressive Learning (SAMPL) method.
-Specifically, we utilize SAM (Kirillov et al.
-2023) to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying size ratios.
+Specifically, we utilize SAM Kirillov et al.
+(2023) to identify HOI-relevant regions and subsequently apply a dynamic masking strategy with varying size ratios.
 This approach enables the model to focus on both holistic and partial information, thereby enhancing its robustness throughout the optimization process.
@@
   • SAMPL yields SOTA in terms of robustness, surpassing prior HOI models under both clean and corrupted conditions.
-Related Work
-Human Object Interaction Detection.
-Human-object interaction (HOI) detection (Yao and Fei-Fei 2010; Fang et al.
-2018; Gupta, Schwing, and Hoiem 2019; Li et al.
-2019; Wan et al.
-2019; Park et al.
-2018; Xu et al.
-2020; Yu, Liu, and Odobez 2018; Zhang et al.
-2015; Zhou et al.
-2020) aims to automatically identify and interpret relationships between humans and objects, typically divided into two categories: two-stage and one-stage approaches.
-Two-stage HOI detection methods (Qi et al.
-2018; Xu et al.
-2019; Liang et al.
-2021; Zhou and Chi 2019) first detect humans and objects independently before classifying interactions.
-Early CNN-based approaches (Xu et al.
-2019; Zhou and Chi 2019; Qi et al.
-2018) mainly utilized spatial contexts for interaction modeling.
-Later graph-based methods (Zhou and Chi 2019; Xu et al.
-2019; Ulutan, Iftekhar, and Manjunath 2020; Wang et al.
-2021; Zhang, Campbell, and Gould 2021) improved relational reasoning explicitly.
-Subsequent methods further enhanced precision via pose-aware representations (Wan et al.
-2019), transformer-based unpaired modeling (Zhang, Campbell, and Gould 2022), and integration-decomposition strategies (Li et al.
-2020).
-One-stage HOI detection methods (Yuan et al.
-2023; Zhang et al.
-2023) achieve efficiency by directly predicting human-object interactions end-to-end.
-Early methods exploited parallel point detection (Liao et al.
-2020) and union-level modeling (Kim et al.
-2020), while subsequent approaches refined interaction localization via adaptive glance-and-gaze mechanisms (Zhong et al.
-2021).
-Transformer-based methods enhanced relational context modeling (Kim, Jung, and Cho 2023) and language-image fusion (RLIPv2)(Yuan et al.
-2023), significantly improving performance.
-Recently, vision-language models (VLMs) were leveraged to boost generalization across unseen interactions (VLM-HOI(Kang et al.
-2024)) and zero-shot triplet predictions (EZ-HOI (Lei, Wang, and Tan 2024)).
-Robustness-specific methods further integrated uncertainty estimation (UAHOI (Chen, Chen, and Yang 2024)) to improve stability under perturbations.
-Nevertheless, all these approaches primarily evaluate performance on clean data, neglecting real-world corruptions.
-To address this gap, we introduce RoHOI, the first HOI robustness benchmark encompassing $20$ test-time corruptions.
+Figure 1: Our RoHOI dataset comprises 20 types of algorithmically generated corruptions, systematically categorized into four groups.
+Each corruption type has five levels of severity, leading to a total of 100 distinct corruptions.
+These corruptions simulate realistic semantic and structural disturbances specifically encountered in practical applications.
+2 Related Work
+Human-Object Interaction Detection.
+Human-object interaction (HOI) detection Yao & Fei-Fei (2010); Fang et al.
+(2018); Gupta et al.
+(2019); Li et al.
+(2019); Wan et al.
+(2019); Park et al.
+(2018); Xu et al.
+(2020); Yu et al.
+(2018); Zhang et al.
+(2015); Zhou et al.
+(2020) aims to automatically identify and interpret relationships between humans and objects, typically divided into two categories: two-stage and one-stage approaches.
+Two-stage HOI detection methods Qi et al.
+(2018); Xu et al.
+(2019); Liang et al.
+(2021); Zhou & Chi (2019) first detect humans and objects independently before classifying interactions.
+Early CNN-based approaches Xu et al.
+(2019); Zhou & Chi (2019); Qi et al.
+(2018) mainly utilized spatial contexts for interaction modeling.
+Later graph-based methods Zhou & Chi (2019); Xu et al.
+(2019); Ulutan et al.
+(2020); Wang et al.
+(2021); Zhang et al.
+(2021b) improved relational reasoning explicitly.
+Subsequent methods further enhanced precision via pose-aware representations Wan et al.
+(2019), transformer-based unpaired modeling Zhang et al.
+(2022a), and integration-decomposition strategies Li et al.
+(2020).
+One-stage HOI detection methods Yuan et al.
+(2023); Zhang et al.
+(2023) predict human–object interactions in a single forward pass, improving efficiency.
+Early approaches rely on point-level matching (e.g.
+PPDM Liao et al.
+(2020)) or union-level ROI features (e.g.
+UnionDet Kim et al.
+(2020)).
+Later methods refine localization via glance-and-gaze attention (e.g.
+Zhong et al.
+(2021)).
+Transformer-based pipelines further enrich relational context, and methods like RLIPv2 Yuan et al.
+(2023) combine vision and language modalities for enhanced semantic generalization.
+Recent trends explore generative priors and robust-aware training.
+DIFFUSIONHOI Li et al.
+(2024) integrates a text-to-image diffusion model to derive relation-aware prompts, while HOI-IDiff Hui et al.
+(2025) recasts each human–object prediction as an “HOI image” and generates it with a customized diffusion process.
+UAHOI Chen et al.
+(2024a) explicitly models prediction uncertainty and incorporates it into the objective, yielding more stable interaction predictions.
+Although promising, these approaches are still evaluated on clean benchmarks.
+In contrast, our work introduces RoHOI, a dedicated benchmark of $20$ corruption types, and proposes SAMPL to strengthen robustness under perturbation.
 Visual Robustness Benchmarks.
-Robustness benchmarks have been widely established to assess visual model resilience across various tasks (Chattopadhyay et al.
-2021; Wang et al.
-2024; Sarkar et al.
-2024; Schiappa et al.
-2024; An et al.
-2024; Liu et al.
-2024b; Yi et al.
-2024; Kamann and Rother 2020; Ma et al.
-2024b; Hao et al.
-2024; Xie et al.
-2025; Hao et al.
-2025; Liu et al.
-2024c; Michaelis et al.
-2019; Zeng et al.
-2024), including image classification (ARES-Bench (Liu et al.
-2024a)), visual question answering (Ishmam et al.
-2024), common corruption evaluation (ImageNet-C (Hendrycks and Dietterich 2019)), attribute editing robustness (ImageNet-E (Li et al.
-2023)), one-shot skeleton-based action recognition under occlusion (Peng et al.
-2023), and robust document layout analysis (RoDLA (Chen et al.
-2024b)).
+Robustness benchmarks have been widely established to assess visual model resilience across various tasks Chattopadhyay et al.
+(2021); Wang et al.
+(2024); Sarkar et al.
+(2024); Schiappa et al.
+(2024); An et al.
+(2024); Liu et al.
+(2024b); Yi et al.
+(2024); Kamann & Rother (2020); Ma et al.
+(2024b); Hao et al.
+(2024); Xie et al.
+(2025); Hao et al.
+(2025); Liu et al.
+(2024c); Michaelis et al.
+(2019); Zeng et al.
+(2024), including image classification (ARES-Bench Liu et al.
+(2024a)), visual question answering Ishmam et al.
+(2024), common corruption evaluation (ImageNet-C Hendrycks & Dietterich (2019)), attribute editing robustness (ImageNet-E Li et al.
+(2023)), one-shot skeleton-based action recognition under occlusion Peng et al.
+(2023), and robust document layout analysis (RoDLA Chen et al.
+(2024c)).
 However, robustness benchmarks explicitly designed for HOI detection have not been explored.
 In this work, we address the uncharted robust HOI detection area by constructing the new RoHOI benchmark and by putting forward the SAMPL method to achieve more reliable HOI detection performance.
-Benchmark
+3 Benchmark
 Corruptions.
 To rigorously assess the robustness of HOI detection, we introduce our dedicated RoHOI benchmark, designed to evaluate model performance under real-world corruptions.
-Existing benchmarks such as ImageNet-C (Hendrycks and Dietterich 2019) primarily cover low-level pixel disturbances like brightness or blur.
+Existing benchmarks such as ImageNet-C Hendrycks & Dietterich (2019) primarily cover low-level pixel disturbances like brightness or blur.
 In contrast, RoHOI explicitly includes semantic-level corruptions specifically tailored to challenge spatial and relational reasoning critical for HOI detection.
@@
 To evaluate robustness under real-world corruptions, we use two widely adopted HOI detection datasets: HICO-DET and V-COCO.
-HICO-DET (Wang et al.
-2020) contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories.
+HICO-DET Wang et al.
+(2020) contains $47,776$ images with $600$ interaction classes derived from $117$ action classes and $80$ object categories.
 Each image is annotated with human-object pairs and their corresponding interactions, making it a benchmark for evaluating diverse HOI scenarios.
-V-COCO (Gupta and Malik 2015) is a subset of MS-COCO (Lin et al.
-2014) with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes.
+V-COCO Gupta & Malik (2015) is a subset of MS-COCO Lin et al.
+(2014) with $10,346$ images, $16,199$ human instances, $81$ object categories and $29$ verb classes.
 Unlike HICO-DET, which provides a fine-grained benchmark, V-COCO evaluates model performance on common human-object interactions.
@@
 Baselines.
-To systematically evaluate the robustness of HOI detection models, we benchmark both two-stage and one-stage architectures, including transformer-based and vision-language pretraining approaches.
-Two-stage methods first detect objects before predicting interactions.
-We evaluate QPIC (Tamura, Ohashi, and Yoshinaga 2021) and UPT (Zhang, Campbell, and Gould 2022), both of which use object detectors to extract human-object proposals before interaction classification.
-We include CDN (Zhang et al.
-2021), which refines interaction reasoning via a cascade structure, and DiffHOI (Yang et al.
-2023), which leverages text-to-image diffusion priors to improve rare interaction recognition.
-One-stage methods directly predict interactions without explicit object detection, improving efficiency.
-We evaluate FGAHOI (Ma et al.
-2024a), which enhances spatial alignment using fine-grained anchors, and QAHOI (Chen and Yanai 2023), which integrates query-based attention for interaction localization.
-Additionally, we include MUREN (Kim, Jung, and Cho 2023), which refines relational context modeling, and SOV-STG (Chen, Wang, and Yanai 2023), which optimizes decoder attention for verb-object association.
-GEN-VLKT (Liao et al.
-2022) incorporates visual-language pretraining to improve interaction recognition.
-RLIPv2 (Yuan et al.
-2023) incorporates region-language alignment for fine-grained interaction reasoning.
+To systematically assess robustness, we benchmark representative two-stage and one-stage HOI detectors spanning transformer pipelines and vision–language training paradigms.
+For the two-stage family, we include UPT Zhang et al.
+(2022a), which detects humans and objects before interaction classification using a unary–pairwise transformer, and GEN-VLKT Liao et al.
+(2022), which adopts a guided-embedding association design and a visual–linguistic knowledge transfer (VLKT) strategy that transfers CLIP text embeddings via initialization and a mimic loss rather than performing generic VLM pretraining.
+For the one-stage family, we evaluate QPIC Tamura et al.
+(2021) as a query-based pairwise transformer detector, CDN Zhang et al.
+(2021a) as an end-to-end cascade disentangling network, and DiffHOI Yang et al.
+(2023) which injects text-to-image diffusion priors to enrich relational cues.
+We further include FGAHOI Ma et al.
+(2024a) with fine-grained anchors, QAHOI Chen & Yanai (2023) with query-based anchors, MUREN Kim et al.
+(2023) with multiplex relational context learning, and SOV-STG Chen et al.
+(2023) with subject–object–verb decoding guided by specific-target denoising.
+Finally, we consider RLIPv2 Yuan et al.
+(2023), which incorporates region-language alignment for fine-grained interaction reasoning.
 Additionally, we develop SAMPL, which is built upon RLIPv2 and incorporates semantic-aware masking-based progressive learning.
@@
 Evaluation Metrics.
-We evaluate the clean test set of HICO-DET (Wang et al.
-2020) using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare (<<10 instances), and non-rare ($\geq$10 instances) categories (Yuan et al.
-2023).
-For the clean test set of V-COCO (Gupta and Malik 2015), we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$.
+We evaluate the clean test set of HICO-DET Wang et al.
+(2020) using mAP at Intersection over Union (IoU) $0.5$ for all (Full), rare (<<10 instances), and non-rare ($\geq$10 instances) categories Yuan et al.
+(2023).
+For the clean test set of V-COCO Gupta & Malik (2015), we use $\mathrm{AP}_{\text{role}}$ for interaction detection, with $\mathrm{AP}^{\#1}_{\text{role}}$ for action recognition and $\mathrm{AP}^{\#2}_{\text{role}}$ requiring object localization at IoU $0.5$.
 To evaluate HOI detection performance on our robustness benchmark, we introduce two robustness-focused metrics: Mean Robustness Index (MRI) and Composite Robustness Index (CRI).
-MRI, following the definition in prior work (Hendrycks and Dietterich 2019), measures average performance under corruptions but overlooks stability across severity levels.
+MRI, following the definition in prior work Hendrycks & Dietterich (2019), measures average performance under corruptions but overlooks stability across severity levels.
 To address this, we propose CRI, which penalizes performance variance while normalizing against clean-set accuracy, ensuring models with consistent robustness are appropriately rewarded.
 Mean Robustness Index (MRI) serves as a baseline measure of a model’s average performance across corruption types and severities:
-$\text{MRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{1}{L_{c}}\sum_{l=1}^{L_{c}}mAP_{c,l}\right),$ (1)
-where $C$ denotes the number of corruption types, $L_{c}$ represents the number of severity levels for the $c$-th corruption, and $mAP_{c,l}$ is the model’s performance at level $l$ of corruption $c$.
-While MRI provides an overall robustness estimate, it does not account for performance fluctuations across corruption levels.
-Composite Robustness Index (CRI) addresses MRI’s limitations. It penalizes performance instability across corruption levels while normalizing against clean-set performance:
-$\text{CRI}=\frac{1}{C}\sum_{c=1}^{C}\left(\frac{\overline{mAP}_{c}}{mAP_{clean}}\cdot\frac{1}{\log(1+\sigma_{c})+1}\right),$ (2)
-where $\overline{mAP}_{c}$ is the mean mAP for corruption type $c$, and $\sigma_{c}$ is the standard deviation across corruption levels.
-The clean-set performance $mAP_{clean}$ is defined as mAP (Full) for HICO-DET and $\mathrm{AP}^{\#2}_{\text{role}}$ for V-COCO, where $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5.
-The penalty term $\frac{1}{\log(1+\sigma_{c})+1}$ reduces scores for models with high variance, ensuring that robustness accounts for both overall performance and stability.
-By integrating absolute accuracy with degradation consistency, CRI favors models that maintain stable performance across corruption levels, mitigating the impact of abrupt performance drops.
-For completeness, we also report standard HICO-DET mAP and V-COCO role-based AP scores, following prior benchmarks (detailed in Supplementary).
-Method
+$\mathrm{MRI}\;=\;\frac{1}{C}\sum_{c=1}^{C}\left(\frac{1}{L_{c}}\sum_{l=1}^{L_{c}}M_{c,l}\right).$ (1)
+While MRI provides an overall robustness estimate, it does not capture variability across severities.
+Composite Robustness Index (CRI) counterbalances this by penalizing instability and normalizing to clean-set performance:
+$\mathrm{CRI}\;=\;\frac{1}{C}\sum_{c=1}^{C}\left(\frac{\overline{M}_{c}}{M_{\text{clean}}}\cdot\frac{1}{\log\!\big(1+\sigma_{c}\big)+1}\right).$ (2)
+Conventions for MRI/CRI.
+$C$ is the number of corruption types; $L_{c}$ is the number of severity levels for corruption $c$.
+$M_{c,l}$ denotes the task metric evaluated at level $l$ of corruption $c$; $\overline{M}_{c}$ and $\sigma_{c}$ are, respectively, the mean and standard deviation of $\{M_{c,l}\}_{l=1}^{L_{c}}$; $M_{\text{clean}}$ is the clean-set score.
+Unless otherwise specified, we instantiate $M$ as mAP (Full) on HICO-DET and $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO.
+By combining absolute accuracy with stability across severities, CRI favors models that retain consistent performance under progressively challenging corruptions, complementing the mean-focused view provided by MRI.
+4 Method
 In this section, we present our novel Semantic-Aware Masking-based Progressive Learning (SAMPL), aiming to enhance model robustness by progressively guiding models from holistic feature reliance towards robust contextual inference.
-Unlike standard augmentation methods, SAMPL employs adaptive semantic masking guided by SAM (Kirillov et al.
-2023), dynamically adjusting masking severity based on validation performance.
+Unlike standard augmentation methods, SAMPL employs adaptive semantic masking guided by SAM Kirillov et al.
+(2023), dynamically adjusting masking severity based on validation performance.
 This structured semantic-level masking strategy explicitly promotes robust contextual reasoning by training models to infer interactions from progressively restricted semantic regions, thus effectively improving general robustness against a variety of realistic disturbances.
@@
 11:     $N(\omega_{p}):=N(\omega_{p})+1$. $\Delta Q_{\max}\leftarrow$-$\infty$
-12:   else  Train model on $\omega^{*}$. $N(\omega_{p}):=N(\omega_{p})+1$.
+12:   else  Train model on $\omega^{*}$. $N(\omega^{*}):=N(\omega^{*})+1$.
 13:   end if
@@
 16:end for
+Table 1: Comparison of HOI detection methods on HICO-DET and V-COCO under clean and corrupted settings.
+Results on HICO-DET are reported on Full, Rare and Non-Rare splits; V-COCO results use $\mathrm{AP}^{\#1}_{\text{role}}$ and $\mathrm{AP}^{\#2}_{\text{role}}$.
+Columns Backb., Det., VLM indicate whether a model uses a pretrained backbone, a detection-model pretraining, or vision–language pretraining.
+MRI and CRI quantify robustness under corruptions (higher is better).
 […]
@@
 […]
+One-Stage methods
 […]
 […]
-[…]
-One-Stage methods
 […]
@@
 The whole procedure is summarized in Alg. 1.
-Experimental Results
+5 Experimental Results
 Implementation Details.
-Our SAMPL method is built upon RLIPv2 (Yuan et al.
-2023) with a Swin-Tiny (Liu et al.
-2021) backbone and has $214M$ trainable parameters.
-The backbone is pretrained on Visual Genome (Krishna et al.
-2017), MS-COCO (Lin et al.
-2014) and Object365 (Shao et al.
-2019), and fine-tuned on the HICO-DET or V-COCO datasets.
-We use the AdamW optimizer (Loshchilov and Hutter 2019) with an initial learning rate of 1ee-4 (backbone and text encoder at 1ee-5), linear warmup followed by cosine decay, weight decay of 1ee-4, and batch size $2$ per GPU with gradient accumulation over $2$ steps.
+Our SAMPL method is built upon RLIPv2 Yuan et al.
+(2023) with a Swin-Tiny Liu et al.
+(2021) backbone and has $214M$ trainable parameters.
+The backbone is pretrained on Visual Genome Krishna et al.
+(2017), MS-COCO Lin et al.
+(2014) and Object365 Shao et al.
+(2019), and fine-tuned on the HICO-DET or V-COCO datasets.
+We use the AdamW optimizer Loshchilov & Hutter (2019) with an initial learning rate of 1ee-4 (backbone and text encoder at 1ee-5), linear warmup followed by cosine decay, weight decay of 1ee-4, and batch size $2$ per GPU with gradient accumulation over $2$ steps.
 Models are trained for $80$ epochs using cross-entropy loss for object classification, focal loss for verb classification, and L1 + GIoU loss for bounding, with KL divergence for latent space regularization.
@@
 For the dynamic severity adjustment mechanism, we set the initial threshold as $\tau_{\text{init}}=0.15$ and $\epsilon=10^{-6}$.
-Figure 2: Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks.
+(a) Refer to caption
+(b) Refer to caption
+Figure 2: (a) Performance degradation of RLIPv2 Yuan et al.
+(2023) on HICO-DET and V-COCO under RoHOI.
+The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO.
+Clean dataset performance serves as a reference, with shaded areas indicating corruption impact.
+(b) Comparison of SAMPL with strong two-stage and one-stage baselines trained on V-COCO under RoHOI benchmarks.
 The left bar chart groups performance by corruption categories, while the right radar chart compares models across five severity levels for three corruptions.
@@
 Details are shown in Sec.
-Benchmark.
+3 and Appendix.
 Analysis of the Benchmark.
-Table 1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET (Wang et al.
-2020) and V-COCO (Gupta and Malik 2015) datasets.
-SAMPL surpasses RLIPv2 (Yuan et al.
-2023), improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations.
+Tab.
+1 presents the performance comparison between SAMPL and existing one-stage and two-stage HOI detection methods on HICO-DET Wang et al.
+(2020) and V-COCO Gupta & Malik (2015) datasets.
+SAMPL surpasses RLIPv2 Yuan et al.
+(2023), improving MRI by $3.78$ and CRI by $12.5\%$ on V-COCO, demonstrating superior robustness against visual degradations.
 On HICO-DET, SAMPL achieves a $3.6\%$ higher CRI while maintaining comparable MRI, ensuring strong generalization without compromising clean-set accuracy.
 These results are further supported by Fig.
-2, where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the supplementary).
+2 (b), where SAMPL consistently outperforms two-stage and one-stage baselines across all $4$ corruption categories in the RoHOI benchmark, achieving the highest MRI in each group and demonstrating improvements over the current SOTA model at every severity level (detailed in the Appendix).
 While extreme noise levels still pose challenges, SAMPL significantly improves robustness compared to prior methods by progressively training the model with semantic-aware occlusions, enabling it to learn context-driven features that are resilient to diverse and severe corruptions.
@@
 Through a comprehensive analysis of representative HOI models, we derive the following key findings.
-Observation 1: Large-scale vision-language pretraining significantly enhances the robustness of transformer-based models.
-Experiments on the clean and corrupted test set (Table 1) reveal that transformer-based architectures exhibit robustness variations dependent on their pretraining strategies.
-RLIPv2, which leverages large-scale multimodal pretraining, achieves substantially higher robustness scores in contrast to GEN-VLKT, SOV-STG and other transformer-based methods.
-This suggests that large-scale vision-language pretraining provides a more structured relational feature space, mitigating feature drift and localization sensitivity in HOI detection tasks.
-Furthermore, two-stage models that utilize pretrained detector backbone, e.g., UPT, show only moderate robustness improvements, highlighting the importance of diverse relational supervision beyond object-centric datasets.
-Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 (Yuan et al.
-2023).
+Observation 1: Vision–language relational pretraining is associated with higher corruption-robustness.
+Within each dataset (Tab.
+1), HOI detectors trained with relational vision–language objectives occupy the top tier of MRI/CRI, while transformer-based counterparts without such cross-modal relational supervision lag behind.
+A plausible mechanism is that relational fusion during pretraining aligns verb–object evidence and regularizes interaction features against distribution shift; We interpret this pattern as associative rather than causal; backbone capacity remains a potential confounder.
+Figure 3: Qualitative comparison of HOI detection under corruption between SAMPL (ours) and RLIPv2 Yuan et al.
+(2023).
 The first row shows uncorrupted images, while the second and third rows present SAMPL and RLIPv2 results on corrupted images.
 Missed predictions are in dashed red, false interactions in red, and correct ones in blue.
-Observation 2: Object localization strategies dictate robustness in two-stage architectures.
-Two-stage methods show varying robustness across different object localization strategies.
-Region-proposal-based models (Zhang et al.
-2021) achieve high clean-set accuracy but degrade under corruption due to error propagation in localization.
-In contrast, query-based methods (Zhang, Campbell, and Gould 2022) adapt object representations dynamically, enhancing robustness on HICO-DET.
-However, on V-COCO, UPT with a larger backbone surprisingly shows lower CRI than CDN with a smaller backbone, likely attributed to V-COCO’s limited interaction diversity.
-This suggests that query-based detection benefits from diverse datasets but struggles with constrained interaction distributions.
+[…]
+[…]
+Observation 4: Progressive semantic masking primarily improves stability (CRI) and smooths severity–performance curves.
+Relative to a strong vision–language pretrained baseline in Tab.
+1, SAMPL consistently increases CRI within each dataset, while MRI gains are moderate or dataset-dependent.
+This pattern aligns with evidence that self-attentive encoders are sensitive to patch-level corruptions: structure-aware, progressive masking encourages attention to retain relational evidence under partial visibility, yielding smoother severity–performance profiles without sacrificing clean accuracy.
+Table 2: Ablation studies on V-COCO.
+(a) Ablation on various perturbation strategies during training: SAM-guided HOI-aware Semantic Masking (SAMSM), Adaptive Gaussian Noise Perturbation (AGNP), and Randomized Pixelation Perturbation (RPP).
+(b) Ablation on training with different severity level combinations for SAMSM: each instance is masked based on 40%–80% of its bounding box height and width, using three candidate levels during training.
+(a) Training perturbation strategies
 […]
@@
 […]
-Table 2: Ablation study on various perturbation strategies during training.
-We evaluate SAM-guided HOI-aware Semantic Masking (SAMSM), Adaptive Gaussian Noise Perturbation (AGNP), and Randomized Pixelation Perturbation (RPP) on the V-COCO.
-Observation 3: Interaction representation entanglement amplifies robustness degradation under perturbations.
-Models with tightly coupled interaction representations achieve higher clean-set accuracy but are more prone to corruption due to feature distortion propagation.
-MUREN, with multi-level relational context learning, suffers the largest MRI and CRI drop, showing that excessive entanglement weakens robustness.
-Similarly, GEN-VLKT-L’s global self-attention leads to greater degradation than CDN-L’s decoupled interaction classification, while QAHOI’s adaptive query-based approach improves robustness by dynamically refining predictions.
-This underscores a key trade-off: stronger feature coupling boosts accuracy but reduces resilience to corruptions.
-Observation 4: Dataset interaction diversity directly influences robustness generalization.
-Models trained on HIO-DET achieve higher MRI and CRI than V-COCO-trained models due to greater interaction diversity, enhancing resilience.
-However, SAMPL shows stronger robustness gains on V-COCO, as its progressive learning strategy compensates for dataset limitations, achieving comparable robustness with less data and fewer iterations.
-This underscores the importance of diverse interactions for general robustness and adaptive training for faster improvements in smaller datasets.
-Structural design considerations in both architectures and training paradigms are crucial for balancing recognition accuracy and robustness in real-world applications.
+(b) SAMSM cover ratio combinations
 […]
@@
 […]
-Table 3: Ablation study on training with different severity level combinations.
-SAMSM masks each instance based on 40%–80% of its bounding box height and width, using three candidate levels during training.
 Analysis of the Ablation Study.
 To investigate the impact of perturbation-based training on model robustness, we conduct a series of ablation experiments on the V-COCO dataset.
-The results, presented in Tables 2 and 3, analyze the effects of different perturbation types and perturbation level settings on model performance.
-Ablation of the Semantic-Aware Masking: Table 2 evaluates three perturbation strategies applied independently during training, excluding our progressive learning method to isolate their direct impact.
-SAM-guided HOI aware Semantic Masking (SAMSM) is applied in 50% of training epochs, occluding instance’s regions while preserving contextual cues.
+The results, presented in Tab.
+2, analyze the effects of different perturbation types and perturbation level settings on model performance.
+Ablation of the Semantic-Aware Masking: Tab.
+2 (a) evaluates three perturbation strategies applied independently during training, excluding our progressive learning method to isolate their direct impact.
+SAM guided HOI-aware Semantic Masking (SAMSM) is applied in 50% of training epochs, occluding instances’ regions while preserving contextual cues.
 Adaptive Gaussian Noise Perturbation (AGNP) introduces feature-level noise with a 50% probability, while Randomized Pixelation Perturbation (RPP) pixelates input images with the same probability as data augmentation.
@@
 Despite expectations that RPP would improve robustness—given that pixelation causes the most significant performance drop in baseline corruption tests (as shown in Fig.
-4)—its randomized nature indiscriminately distorts both essential and redundant information, making adaptation less effective.
+2 (a))—its randomized nature indiscriminately distorts both essential and redundant information, making adaptation less effective.
 These results suggest that robustness in HOI detection benefits more from structured perturbations that challenge object-awareness rather than indiscriminate degradations that introduce confounding artifacts.
-Figure 4: Performance degradation of RLIPv2 (Yuan et al.
-2023) on HICO-DET and V-COCO under RoHOI.
-The left radar chart shows the relative mAP (Full) drop on HICO-DET, while the right shows the decline in $\mathrm{AP}^{\#2}_{\text{role}}$ on V-COCO.
-Clean dataset performance serves as a reference, with shaded areas indicating corruption impact.
-Effect of Perturbation Levels: Table 3 analyzes the impact of masking severity on robustness.
+Effect of Perturbation Levels: Tab.
+2 (b) analyzes the impact of masking severity on robustness.
 SAMSM occludes $40\%$$80\%$ of each instance’s bounding box, with the $40\%/50\%/60\%$ setting achieving the best trade-off, maximizing MRI and CRI while maintaining competitive AP scores.
@@
 However, both models remain vulnerable under severe degradations, where dataset priors dominate predictions.
-Conclusion
-In this paper, we introduced RoHOI, the first robustness benchmark for HOI detection, evaluating models under $20$ real-world perturbations.
-Our analysis shows that existing models suffer significant degradation, highlighting the need for more robust approaches.
-To address this, we propose SAMPL, a Semantic-Aware Masking-based Progressive Learning method that utilizes SAM-generated region-aware masks and dynamically adjusts masking severity.
-SAMPL enhances robustness while maintaining high accuracy on clean datasets, outperforming state-of-the-art methods.
-By bridging the gap between clean-data performance and real-world robustness, our work advances reliable and adaptable HOI detection models for practical use.
-Appendix A Social Impact and Limitations
+6 Conclusion
+In this work, we introduced RoHOI, the first benchmark dedicated to evaluating the robustness of Human-Object Interaction (HOI) detection models under 20 diverse real-world corruptions.
+Our analysis revealed that even state-of-the-art approaches degrade substantially under perturbations, underscoring the need for robustness-focused evaluation.
+To address this challenge, we proposed SAMPL (Semantic-Aware Masking-based Progressive Learning), which leverages SAM-generated region-aware masks and a dynamic progressive masking strategy to strengthen resilience.
+Extensive experiments on HICO-DET and V-COCO showed that SAMPL consistently outperforms strong baselines under corruption while preserving state-of-the-art performance on clean benchmarks.
+Beyond performance gains, our results highlight that clean accuracy and robustness are only partially coupled, that structured perturbations provide more effective robustness cues than random noise, and that architecture design plays a crucial role in stability.
+While RoHOI captures a broad range of corruptions, it still focuses on synthetic perturbations; future work should explore robustness under real-world domain shifts, extreme degradations, and fairness-aware evaluation.
+We envision RoHOI as a catalyst for advancing robustness research in HOI detection, and SAMPL as an initial step toward building reliable and deployable models for safety-critical applications.
+Acknowledgements
+The project is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – SFB 1574 – 471687386.
+This work was supported in part by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026), the MWK through the Cooperative Graduate School Accessibility through AI-based Assistive Technology (KATE) under Grant BW6-03, and in part by the BMBF through a fellowship within the IFI program of the German Academic Exchange Service (DAAD), in part by the HoreKA@KIT supercomputer partition, in part by the National Natural Science Foundation of China (No.
+62473139), in part by the Hunan Provincial Research and Development Project (Grant No.
+2025QK3019), and in part by the Open Research Project of the State Key Laboratory of Industrial Control Technology, China (Grant No.
+ICT2025B20).
+Appendix A Appendix
+A.1 Social Impact and Limitations
 Social Impact: We introduce RoHOI, the first benchmark for assessing HOI detection robustness under real-world corruptions, incorporating 20 degradation types to bridge the gap between theoretical advancements and deployment.
@@
 While SAMPL improves robustness, it still struggles with severe distortions (e.g., moire patterns, glass blur), indicating the need for further architectural advancements.
-Appendix B Standard HOI Evaluation Metrics
-This section provides a detailed explanation of the evaluation metrics used for HICO-DET (Xu et al.
-2019) and V-COCO (Gupta and Malik 2015) in our main paper.
+A.2 Standard HOI Evaluation Metrics
+This section provides a detailed explanation of the evaluation metrics used for HICO-DET Xu et al.
+(2019) and V-COCO Gupta & Malik (2015) in our main paper.
 These include the standard mAP metrics for HICO-DET and AProle variations for V-COCO.
-$\bm{\mathrm{mAP}}\,\textbf{(Full, Rare, Non-Rare)}$
+A.2.1 $\mathbf{mAP}$: Full, Rare, Non–Rare
 We evaluate models on HICO-DET using the standard mAP under an IoU threshold of 0.5.
 The reported metrics include mAP (Full), which evaluates all 600 HOI categories; mAP (Rare), assessing performance on low-frequency categories with fewer than 10 instances; and mAP (Non-Rare), focusing on categories with at least 10 instances.
-$\bm{\mathrm{AP}^{\#1}_{\text{role}}}$ and $\bm{\mathrm{AP}^{\#2}_{\text{role}}}$
+A.2.2 $\mathbf{AP}^{\#1}_{\text{role}}\ \text{and}\ \mathbf{AP}^{\#2}_{\text{role}}$
 We evaluate V-COCO using $\mathrm{AP}_{\text{role}}$, which measures interaction detection accuracy.
@@
 $\mathrm{AP}^{\#2}_{\text{role}}$ requires both action and object localization, with an IoU threshold of 0.5.
-Model MB DB GauB GB GauN ShN S&P JPEG SN PL EXP RE OCC VE MP SC ET PD PIX ZB
-CDN-L (Zhang et al. 2021) 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
-GEN-VLKT-L (Liao et al. 2022) 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
-MUREN (Kim, Jung, and Cho 2023) 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
-RLIPv2-ParSeDA (Yuan et al. 2023) 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
-[…]
-Table 4: Comparison of MRI scores between our SAMPL method and the best-performing one-stage and two-stage HOI detection models on the RoHOI benchmark.
-The table presents results across 20 corruption types, evaluating model robustness under diverse corruptions.
-Corruption abbreviations follow the definitions provided in the main paper.
-Level 1 Level 2 Level 3 Level 4 Level 5
-Motion Blur (MB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Defocus Blur (DB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Gaussian Blur (GauB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Glass Blur (GB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 5: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
-Appendix C Corruption Types in the Dataset
+A.3 Corruption Types in the Dataset
 Building upon the general introduction in the main text, this section provides a more detailed breakdown of the $20$ corruption types used in our RoHOI benchmark.
-Optical System-Induced (OS) Artifacts
+A.3.1 Optical System-Induced (OS) Artifacts
 Motion Blur (MB): Blur caused by the relative motion between the camera and the scene during exposure.
@@
 Level 1 Level 2 Level 3 Level 4 Level 5
+Motion Blur (MB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Defocus Blur (DB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Gaussian Blur (GauB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Glass Blur (GB) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 4: Visual examples of Optical System (OS)-Induced artifacts across severity levels.
+Level 1 Level 2 Level 3 Level 4 Level 5
 Salt-and-Pepper Noise (S&P) Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Figure 5: Visual examples of Sensor, Compression and Transmission (SCT) artifacts across severity levels.
-Sensor, Compression and Transmission (SCT) Artifacts
+A.3.2 Sensor, Compression and Transmission (SCT) Artifacts
 Salt-and-Pepper Noise (S&P): Discrete black-and-white pixel noise caused by sensor malfunction or transmission errors.
@@
 (Details shown in Figure 5, row $6$.)
-Environmentally Induced (EI) Artifacts
+A.3.3 Environmentally Induced (EI) Artifacts
 Level 1 Level 2 Level 3 Level 4 Level 5
@@
 (Details shown in Figure 6, row $4$.)
-Geometric and Scene (G&S) Distortions
+A.3.4 Geometric and Scene (G&S) Distortions
 Level 1 Level 2 Level 3 Level 4 Level 5
@@
 (Details shown in Figure 7, row $6$.)
-Purpose of Inclusion
+A.4 Purpose of Inclusion
 These noise types are selected to comprehensively represent real-world scenarios that affect image quality, ranging from optical and sensor limitations to environmental and dynamic scene factors.
 Their inclusion ensures the dataset’s relevance for evaluating model robustness in diverse and realistic conditions.
-Appendix D Benchmark Results on Corruption Types
+A.5 Benchmark Results on Corruption Types
 Table 3 presents the MRI scores of our SAMPL method compared to the strongest one-stage and two-stage HOI detection baselines on the RoHOI benchmark.
@@
 These results validate the effectiveness of our SAMPL strategy in improving model resilience against real-world corruptions.
+Table 3: Comparison of MRI scores between our SAMPL method and the best-performing one-stage and two-stage HOI detection models on the RoHOI benchmark.
+The table presents results across 20 corruption types, evaluating model robustness under diverse corruptions.
+Corruption abbreviations follow the definitions provided in the main paper.
+Model MB DB GauB GB GauN ShN S&P JPEG SN PL EXP RE OCC VE MP SC ET PD PIX ZB
+CDN-L Zhang et al. (2021a) 15.38 10.12 48.73 7.56 41.67 60.92 28.78 56.89 24.24 28.18 64.67 41.06 49.75 67.07 30.67 44.30 37.33 50.79 22.39 22.45
+GEN-VLKT-L Liao et al. (2022) 14.64 9.90 48.27 7.74 40.50 60.63 28.37 56.31 20.40 28.52 64.94 40.48 50.48 67.48 32.42 44.21 38.27 51.15 21.11 23.24
+MUREN Kim et al. (2023) 15.71 12.10 51.08 6.77 41.65 62.59 26.08 56.32 20.46 23.89 66.65 38.55 51.78 69.12 25.63 44.36 38.79 52.34 15.97 22.22
+RLIPv2-ParSeDA Yuan et al. (2023) 25.31 17.84 57.82 12.99 50.91 66.72 50.58 64.68 40.11 38.34 68.39 55.99 54.61 69.82 45.93 59.14 48.14 53.01 26.12 29.73
+[…]